### 20 Dec 2023 v1

# A faster FPRAS for #NFA

Abstract.
Given a non-deterministic finite automaton (NFA) $\mathcal{A}$ with $m$ states, and a natural number $n\in\mathbb{N}$ (presented in unary), the #NFA problem asks to determine the size of the set $\mathcal{L}(\mathcal{A}_{n})$ of words of length $n$ accepted by $\mathcal{A}$. While the corresponding decision problem of checking the emptiness of $\mathcal{L}(\mathcal{A}_{n})$ is solvable in polynomial time, the #NFA problem is known to be #P-hard. Recently, the long-standing open question — whether there is an FPRAS (fully polynomial time randomized approximation scheme) for #NFA — was resolved in (Arenas et al., 2019). The FPRAS due to (Arenas et al., 2019) relies on the interreducibility of counting and sampling, and computes, for each pair of state $q$ and natural number $i\leq n$, a set of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ many uniformly chosen samples from the set of words of length $i$ that have a run ending at $q$. This informative measure — the number of samples maintained per state and length — also affects the overall time complexity with a quadratic dependence. Given the prohibitively high time complexity, in terms of each of the input parameters, of the FPRAS due to (Arenas et al., 2019), and considering the widespread application of approximate counting (and sampling) in various tasks in Computer Science, a natural question arises: Is there a faster FPRAS for #NFA that can pave the way for the practical implementation of approximate #NFA tools? In this work, we demonstrate that significant improvements in time complexity are achievable. Specifically, we have reduced the number of samples required for each state to be independent of $m$, with significantly less dependence on $n$ and $\epsilon$, maintaining only $\widetilde{O}(\frac{n^{4}}{\epsilon^{2}})$ samples per state. We believe our insights may lead to further reductions in time complexity and thus open up a promising avenue for future work towards the practical implementation of tools for approximating #NFA.
The authors decided to forgo the old convention of alphabetical ordering of authors in favor of a randomized ordering, denoted by r⃝. The publicly verifiable record of the randomization is available at https://www.aeaweb.org/journals/policies/random-author-order/search
copyright: noneconference: ; ;
1. Introduction
In this paper, we focus on the following computational problem:
  #NFA:
: Given an NFA $\mathcal{A}=(Q,I,\Delta,F)$ over the binary alphabet $\Sigma=\{0,1\}$ with $m$ states, and a number $n$ in unary, determine $|\mathcal{L}(\mathcal{A}_{n})|$, where $\mathcal{L}(\mathcal{A}_{n})$ is the set of strings of length $n$ that are accepted by $\mathcal{A}$.
The counting problem #NFA and the closely related problem of uniform sampling words of an NFA are fundamental to database applications, and arise when counting and uniformly generating solutions to queries. Beyond database systems, the aforementioned problem finds applications in program verification (Legay et al., 2010), testing of software (Sutton et al., 2007) and distributed systems (Ozkan et al., 2019), probabilistic programming, estimation of information leakage in software (Bang et al., 2016; Gao et al., 2019; Saha et al., 2023), and even music generation (Donzé et al., 2014).
How do we count the number of elements of $\mathcal{L}(\mathcal{A}_{n})$? When $\mathcal{A}$ is a DFA with $m$ states, this question can be solved using a simple dynamic programming algorithm that inductively computes, for every $i\leq n$ and for every state $q$ of $\mathcal{A}$, the number of words of length $i$ that end in state $q$: $\mathsf{count}(q,i)=\sum_{b\in\Sigma}\sum_{p:q=\Delta(p,b)}\mathsf{count}(p,i-1)$ taking overall $O(mn)$ time (assuming $|\Sigma|=O(1)$). This simple algorithm however fails to work when $\mathcal{A}$ is a non-deterministic automaton as the same word maybe accepted using many different runs. Determinizing $\mathcal{A}$ first and then using the same algorithm as for DFAs would give an overall running time of $O(n\cdot 2^{m})$ which is exponential in the size of the description of the input. It turns out that #NFA is, in fact, provably computationally intractable — it is known to be a #P-hard problem (Álvarez and Jenner, 1993).
Given the intractability, and given the widespread applications of this problem, the next natural question then arises — can we approximate #NFA using a polynomial time algorithm? One of the promising paradigms for deriving efficient algorithms for approximating the answer to an intractable problem is the use of randomization to obtain a fully polynomial time randomized approximation scheme (FPRAS). Formally, an FPRAS for a real-valued function $f:\mathcal{D}\to\mathbb{N}$ over some domain $\mathcal{D}$ is a randomized algorithm $A:\mathcal{D}\times(0,1)\times(0,1)\to\mathbb{N}$ that for every input $x\in\mathcal{D}$ and parameters $\epsilon\in(0,1)$ and $\delta\in(0,1)$, runs in time $\textsf{poly}(|x|,\frac{1}{\epsilon},\log(\frac{1}{\delta}))$, and satisfies: ${\mathsf{Pr}}\!\left[\frac{1}{(1+\epsilon)}\cdot f(x)\leq|A(x,\epsilon,\delta)|\leq(1+\epsilon)\cdot f(x)\right]\geq 1-\delta$.
Until very recently, the only approximation algorithms known for #NFA had running times that were quasi-polynomial in the size of the inputs (Gore et al., 1997; Kannan et al., 1995). In a recent breakthrough by Arenas, Croquevielle, Jayram and Riveros (Arenas et al., 2019, 2021), #NFA was finally shown to admit an FPRAS. Even though the problem was shown to admit an FPRAS, the time complexity of the proposed FPRAS is prohibitively large — it takes time $\widetilde{O}\left(m^{17}n^{17}\cdot\frac{1}{\epsilon^{14}}\cdot\log(1/\delta)\right)$, where the notation $\widetilde{O}$ hides the polylogarithmic multiplicative factors in $m$ and $n$. To be applicable in a practical setting like estimating information flow in programs, a polynomial time approximation scheme whose time complexity varies as $O(|P|^{16})$, where $|P|$ is roughly the lines of code, is unlikely to scale even for moderately sized programs. Towards this, in this work, we pursue the goal of desigining a faster algorithm, and propose an FPRAS which runs in time $\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$. The reliance of both FPRASes on the inter-reducibility of sampling and counting necessitates a more informative measure, such as the number of samples maintained for each state. The scheme proposed in  (Arenas et al., 2021) maintains $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ samples per state while the scheme proposed in this work maintains only $O(\frac{n^{4}}{\epsilon^{2}})$ samples, which is independent of $m$ and with a significant reduction in it’s dependence on $n$ and $\epsilon$.
1.1. Overview of our FPRAS
At a high-level, our approach exploits the tight connection between approximate counting and the related problem of (almost) uniform generation. In the context of an NFA $\mathcal{A}$ and a length $n$, the almost uniform generation problem asks for a randomized algorithm (generator) $G$, such that, on each invocation, the probability with which $G$ returns any given string $x\in\mathcal{L}(\mathcal{A}_{n})$ lies in the range $[\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}-\epsilon,\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}+\epsilon]$; here $\epsilon\in(0,1)$ is a parameter. Such a random generator is a polynomial time generator if its running time is polynomial in $|\mathcal{A}|$ and $n$. The existence of a polynomial time almost uniform generator for self-reducible (Jerrum et al., 1986) problems implies the existence of an FPRAS, and vice versa.
Input : NFA $\mathcal{A}$ with $m$ states $Q$, $n\in\mathbb{N}$ (in unary)
1 Unroll the automaton $\mathcal{A}$ into an acyclic graph $\mathcal{A}_{\textsf{unroll}}$ by making $n+1$ copies $\{q^{\ell}\}_{\ell\in[0,n],q\in Q}$ of the states $Q$ and adding transitions between immediate layers foreach $\ell\in\{0,\ldots,n\},q\in Q$ do
2    Compute the estimate $\mathsf{N}({q}^{\ell})$ using prior estimates and samples $\{\mathsf{N}({q}^{i}),\mathsf{S}({q}^{i})\}_{i<\ell,p\in Q}$ Compute a uniform set of samples $\mathsf{S}({q}^{\ell})$ using $\mathsf{N}({q}^{\ell})$ and the prior estimates and samples $\{\mathsf{N}({q}^{i}),\mathsf{S}({q}^{i})\}_{i<\ell,p\in Q}$
return $\mathsf{N}({F}^{\ell})$
Algorithm 1 Algorithm Template: FPRAS for #NFA
This tight connection between counting and sampling was also the high-level idea behind the work of (Arenas et al., 2021). Consequently, our algorithm fits the template procedure of Algorithm 1, also outlined in (Arenas et al., 2021). For technical convenience, the algorithm unrolls the automaton $\mathcal{A}$ into an acyclic graph with $n$ levels; each level $\ell$ containing the $\ell^{\text{th}}$ copy ${q}^{\ell}$ of each state $q$ of $\mathcal{A}$. A key component in the algorithm is to incrementally compute an estimate $\mathsf{N}({q}^{\ell})$ of the size of the set $\mathcal{L}({q}^{\ell})$ of words of length $\ell$ that have a run ending in state $q$.
Next, in order to inductively estimate $\mathsf{N}({q}^{\ell})$, we in turn rely on the estimates of sets $\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ (${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\Sigma=\{0,1\}$); here $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is a ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor state of $q$, i.e., $(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta$ is a transition of $\mathcal{A}$. These estimates can be put to use effectively because:
$\mathcal{L}({q}^{\ell})=\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}\}\cup\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}.$
We remark that $\mathcal{L}({p}^{\ell-1})$ may not be disjoint for different predecessor states $p$. Hence, we cannot estimate $\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ by merely summing the individual estimates $\{\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})\,|\,(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta\}$. To address this, we maintain a set, $\mathsf{S}({p}^{\ell-1})$, for each $p$, which contains samples of strings from $\mathcal{L}({p}^{\ell-1})$. One possible idea is to construct the set $\mathsf{S}({p}^{\ell-1})$ using strings $w$ that have been sampled uniformly at random from $\mathcal{L}({p}^{\ell-1})$ but such a process introduces dependence among the sets $\mathsf{S}({p}^{\ell-1})$ for different $p$. The crucial observation in (Arenas et al., 2021) is that for any $k$, the following is true: for any word $w$ of length $k$, all words from $\mathcal{L}({q}^{\ell})$ with the suffix $w$ can be expressed as $\big{(}\bigcup\limits_{p\in P}\mathcal{L}({p}^{\ell-k})\big{)}\cdot\{w\}$ for some set $P$ of states, where $P$ depends only on $w,q,$ and $k$. This property is referred to as the self-reducible union property. This insight allows us to utilize the self-reducibility property for sampling. We can partition $\mathcal{L}(p^{\ell-1})$ into sets of strings where the last character is either ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}$ or ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}$, and compute estimates for each partition since each partition can be represented as a union of sets for which we (inductively) have access to both samples and estimates. We then sample the final character in proportion to the size of these estimates. In a similar recursive manner, we can sample strings, character by character. That is, we can sample strings by progressively extending suffixes of $w$ backwards! Of course, we must also account for the errors that propagate in the estimates of $\mathsf{N}({q}^{\ell})$.
Our algorithm also relies on the self-reducible union property, and therefore, the basic structure of our algorithm is similar to that of (Arenas et al., 2021) on high-level. There are, however, significant and crucial technical differences that allow us to obtain an FPRAS with significantly lower time complexity. We highlight these differences below. As a first step, we highlight the two key invariants underpinning the algorithm and ensuing analysis in (Arenas et al., 2021):
  1. (1)
    The following condition (Inv-ACJR) holds for every level $\ell$ (here, $\kappa=\frac{nm}{\epsilon}$), with high probability.
    (Inv-ACJR) $\begin{array}[]{l}\xi(\ell)\triangleq\forall q{\in}Q,\forall P{\subseteq}Q,\big{|}\frac{|\mathsf{S}({q}^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathsf{S}({q}^{\ell})|}-\frac{|\mathcal{L}(q^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}\big{|}\leq\frac{1}{\kappa^{3}}\end{array}$
  2. (2)
    $\mathsf{S}({q}^{\ell})$ is close in $(L_{\infty}$ distance)111 In (Arenas et al., 2021), the invariant is stated in a different, but equivalent, formulation: conditioned on $\xi(\ell)$, the set $\mathsf{S}({q}^{\ell})$ follows the distribution of a multi-set obtained by independently sampling every element of $\mathcal{L}(q^{\ell})$ uniformly at random, with replacement. to the distribution of multi-set obtained by independently sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
In contrast, the FPRAS we propose maintains the following two weaker invariants:
  1. (1)
    For every state $q$ and level $\ell$, the event that $\mathsf{N}({q}^{\ell})\in(1\pm\frac{\epsilon}{2n^{2}})|\mathcal{L}(q^{\ell})|$, which we denote as $\texttt{AccurateN}_{q,\ell}$, happens with high probability.
  2. (2)
    The distribution of $\mathsf{S}({q}^{\ell})$ is close, in total variational distance, to the distribution of a multi-set constructed by sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
Two remarks are in order. First, $\xi(\ell)$ implies $\texttt{AccurateN}_{q,\ell}$, and closeness in $L_{\infty}$ is a stringent condition than closeness in total variation distance. Second, the technical arguments of  (Arenas et al., 2021) crucially rely on their invariants and do not work with the aforementioned weaker arguments. Accordingly, this necessitates a significantly different analysis that depends on coupling of the distributions. It’s also noteworthy that, due to our reliance on the weaker condition, our method for estimating $\mathsf{N}({q}^{\ell})$ differs from that in (Arenas et al., 2021). In fact, our approach is based on an adaptation of the classical Monte Carlo-based computation of the union of sets (Karp and Luby, 1985).
For a given state $q$ and subset $P$, achieving a $\frac{1}{\kappa^{3}}$-additive approximation for $\frac{|\mathcal{L}(q^{\ell})\setminus\cup_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}$ with a probability of $1-\eta$ requires, according to standard concentration bounds, that the size of $\mathsf{S}({q}^{\ell})$ be $O(\kappa^{6}\log\eta^{-1})$. To ensure this additive approximation for all states $q$ and subsets $P$, one must consider union bounds over exponentially many events. Specifically, for the analysis to be successful, one must have $\eta^{-1}\in O(2^{mn})$. Consequently, the FPRAS of (Arenas et al., 2021) maintains $O(\kappa^{7})$ samples for each state $q$. In contrast, for $\texttt{AccurateN}_{q,\ell}$ to hold with a probability of $1-\eta$, our FPRAS only maintains $O\left(\frac{n^{4}}{\epsilon^{2}}\right)\log\eta^{-1}$ samples in $\mathsf{S}({q}^{\ell})$. Further, a union bound over only $O(mn)$ events is required, thus setting $\eta^{-1}=O(mn)$ is adequate.
The time complexity of an algorithm that follows the template outlined in Algorithm 1 (including both the FPRAS of (Arenas et al., 2021) as well as the FPRAS developed in this paper), can be expressed as $O(n\cdot(m\cdot n)\cdot(m^{2}n)\cdot|\mathsf{S}({q}^{\ell})|^{2})$ where the first $n$ from the length of every string, the factor $(m\cdot n)$ corresponds to the upper bound on the number of states in the unrolled automaton, and the factor $(m^{2}n)$ corresponds to the time taken for membership check. Accordingly, it is instructive to compare the bound of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ for $|S(q^{\ell})|$ in  (Arenas et al., 2021) to $O(\frac{n^{4}}{\epsilon^{2}})$ in our approach. Remarkably, the number of samples for every state in our approach is independent of $m$, the size of the automaton.
2. Preliminaries
We consider the binary alphabet $\Sigma=\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$. All results in this paper are extendable to alphabets of arbitrary but fixed constant size. A string $w$ over $\Sigma$ is either the empty string $\lambda$ (of length 00) or a non-empty finite sequence $w_{1}w_{2}\ldots w_{k}$ (of length $|w|=k>0$) where each $w_{i}\in\Sigma$. The set of all strings over $\Sigma$ is denoted by $\Sigma^{*}$. A non-deterministic finite automaton is a tuple $\mathcal{A}=(Q,I,\Delta,F)$ where $Q$ is a finite set of states, $I\in Q$ is the initial state, $\Delta\subseteq Q\times\Sigma\times Q$ is the transition relation, and $F$ is the set of accepting states. A run of $w$ on $\mathcal{A}$ is a sequence $\rho=q_{0},q_{1}\ldots,q_{k}$ such that $k=|w|$, $q_{0}=I$ and for every $i<k$, $(q_{i},w_{i+1},q_{i+1})\in\Delta$; $\rho$ is accepting if $q_{k}\in F$. The word $w$ is accepted by $\mathcal{A}$ if there is an accepting run of $w$ on $\mathcal{A}$. The set of strings accepted by $\mathcal{A}$ is $\mathcal{L}(\mathcal{A})$. For a natural number $n\in\mathbb{N}$, the $n^{\text{th}}$ slice of $\mathcal{A}$’s language, $\mathcal{L}(\mathcal{A}_{n})$, is the set of strings of length $n$ in $\mathcal{L}(\mathcal{A})$. For simplicity, this paper assumes a singleton set of accepting states, i.e., $F=\{q_{F}\}$ for some $q_{F}\in Q$. We introduce additional notation for our algorithms. For $q\in Q$ and $b\in\Sigma$, the $b$-predecessors of $q$ are:
$\textsf{Pred}(q,b)=\{p\,|\,(p,b,q)\in\Delta\}.$
For each $q\in Q$ and $1\leq\ell\leq n$, $\mathcal{L}(q^{\ell})$ denotes the set of strings $w$ of length $\ell$ with a run on $\mathcal{A}$ ending in $q$. The rest of the paper focuses on the automaton $\mathcal{A}=(Q,I,\Delta,F)$, and a number $n$ in unary, using $m$ to denote $|Q|$.
3. A Faster FPRAS
Given automaton $\mathcal{A}$ and a number $n$ in unary, we construct the unrolled automaton $\mathcal{A}_{\textsf{unroll}}$. We use the notation ${q}^{\ell}$ to denote the $\ell^{\text{th}}$ copy of state $q\in Q$. We use $\mathcal{L}({q}^{\ell})$ to denote the the set of all distinct strings $w$ of length $\ell$ for which there is a path (labeled by $w$) from the starting state to $q$ (i.e., those strings that go to state ${q}^{\ell}$ in $\mathcal{A}_{\textsf{unroll}}$). We will use $\mathsf{N}({q}^{\ell})$ to denote an estimate of $|\mathcal{L}({q}^{\ell})|$.
In Section 3.1, we formally present our FPRAS for #NFA, together with the statement of its correctness and running time (Theorem 3). Our FPRAS Algorithm 4 calls two subroutines Algorithm 3 and Algorithm 2, each of which are presented next, together with their informal descriptions. In later sections, we present the formal proofs of their correctness.
3.1. Description of the Algorithms
Algorithm 4 presents the main procedure for approximating the size of $\mathcal{L}(\mathcal{A})_{n}$, where $\mathcal{A}$ is the NFA and $n$ is the parameter that is given as input. The algorithm works in a dynamic programming fashion and maintains, for each state $q$ and each length $1\leq\ell\leq n$, a sampled subset $\mathsf{S}({q}^{\ell})$ of $\mathcal{L}({q}^{\ell}))$. As a key step, it uses the sampled sets corresponding to $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}})$ and $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}})$ to construct the desired sampled subset of $\mathcal{L}({q}^{\ell}))$. We first describe the two subroutines the $\texttt{AppUnion}{()}$ and $\mathsf{sample}()$. We start with describing the subroutines first.
1 function AppUnion ${}_{(\epsilon,\delta)}$($\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$)
2     $m\leftarrow\lceil\frac{\sum_{i}\textsf{sz}_{i}}{\max_{i}\textsf{sz}_{i}}\rceil$ $t\leftarrow\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot m}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ $Y\leftarrow 0$ repeat $t$ times:
3          With prob. $p_{i}=\frac{\textsf{sz}_{i}}{\sum_{i}\textsf{sz}_{i}}$, sample $i\in\{1,\ldots,k\}$ if $S_{i}\neq\varnothing$ then
4               $\sigma\leftarrow S_{\ell}\!\boldsymbol{\cdot}\!\mathsf{dequeue}()$
5         else
6               break
7         if $\not\exists j<i,$ such that $O_{j}\!\boldsymbol{\cdot}\!\mathsf{contains}(\sigma)$ then  $Y\leftarrow Y+1$
8    return $\frac{Y}{t}\cdot{\sum_{i}\textsf{sz}_{i}}$
Algorithm 2 Approximating Union of Sets
1 function $\ell$, ${P}^{\ell}$, $w$, $\varphi$,$\beta$, $\eta$)
2     $\eta^{\prime}\leftarrow\eta/4n$ $\beta^{\prime}\leftarrow(1+\beta)^{\ell-1}-1$ if $\ell=0$ then
3          if $\varphi>1$ then return $\bot$ w.p.  ($\varphi$) : return $w$ ||  w.p.  ($1-\varphi$) : return $\bot$
4    else
5          foreach ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$ do
6               ${P}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\bigcup_{p\in{P}^{\ell}}\textsf{Pred}(p,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ // From the above set, remove all states that are not reachable from the initial state.
[…]
8
9         $\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}\leftarrow\frac{\displaystyle\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}{\displaystyle\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}$ w.p.  ($\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$) : ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\leftarrow{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}$ ||  w.p.  ($1-\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$) : ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\leftarrow{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}$ ${P}^{\ell-1}\leftarrow{P}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ $w\leftarrow{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\cdot w$
10    return sample($\ell-1$, ${P}^{\ell-1}$, $w$, $\varphi/\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},\beta,\eta$)
Algorithm 3 Sampling subroutine for $\mathcal{A}_{\textsf{unroll}}$, length $n$
The AppUnion subroutine. AppUnion ( Algorithm 2) takes as input $\epsilon,\delta$ and $\epsilon_{\textsf{sz}}$ and assumes access to sets $T_{1},\dots,T_{k}$. For each set $T_{i}$ the access is given in the form of membership oracle $O_{i}$, a subset $S_{i}$ (obtained by sampling $T_{i}$ with replacements), and an estimate $\textsf{sz}_{i}$ of the size of $T_{i}$. Using these, Algorithm 2 outputs an $(\epsilon,\delta)$ estimate of the size of the union $\cup_{i=1}^{k}T_{i}$.
The algorithm presented herein represents a modification of classical Monte Carlo-based scheme due to Karp and Luby (Karp and Luby, 1985). The proof methodology also shares similarities with  (Karp and Luby, 1985). For an arbitrary $\sigma\in\cup_{i=1}^{k}T_{i}$, let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$. And let set $\mathcal{U}_{1}$ be the set of all the $(\sigma,L(\sigma))$ pairs and let $\mathcal{U}_{2}$ be the set of all $(\sigma,i)$ pairs such that $\sigma\in T_{i}$. Consequently, the cardinality of the set $\mathcal{U}_{1}$ is $|\cup_{i=1}^{k}T_{i}|$, and that of $\mathcal{U}_{2}$ is $\sum_{i=1}^{k}|T_{i}|$. By drawing a sufficient number of samples from $\mathcal{U}_{2}$ and assessing the fraction belonging to $\mathcal{U}_{1}$, one can estimate $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$, which when multiplied by $|\mathcal{U}_{2}|$ provides a reliable estimate for the size of $\mathcal{U}_{1}$.
In the algorithm we take $t$ samples from $\mathcal{U}_{2}$. This involves initially sampling a set $T_{i}$ proportional to the size of the set (in line 2 ) and subsequently uniformly drawing a sample from $T_{i}$ (in line 2). As the exact sizes of the sets $T_{i}$ are unknown , estimates $\textsf{sz}_{i}$ are utilized for proportional sampling (in line 2), introducing an additional factor of error in the final estimate of $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$.This error can be mitigated by adjusting the value of $t$ accordingly.
Upon sampling an $i$ in Line 2, an element is chosen from the sample of the sets $S_{i}$. This mimics a random sample from $T_{i}$, given that $S_{i}$ already contains uniform samples from $T_{i}$. For each sampled index $i$ , a new element from $S_{i}$ is used and removed (Line 2). Initially, $S_{i}$ are sets are guaranteed to have a size of at least thresh, surpassing the expected number of samples required from $T_{i}$ during the $t$ iterations of the loop. If the algorithm ever needs more samples than $|\mathcal{U}_{i}|$ during the $t$ iterations, it counts as an error, but the probability of this occurrence is demonstrated to be very low.
Finally, line 2 verifies whether the sample from $\mathcal{U}_{2}$ belongs to $\mathcal{U}_{1}$. It uses the oracles to the sets to check this. The variable $Y$ tallies the number of samples from $\mathcal{U}_{1}$, and after $t$ iterations, $Y/t$ provides a estimation for $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$. The final output is the product of this value with $(\sum_{i}\textsf{sz}_{i})$, the estimate for $|\mathcal{U}_{2}|$.
Theorem 1 characterizes the formal guarantee of Algorithm 2.
Theorem 1.
Let $\epsilon,\delta>0$, and let $\Omega$ be some set. Let $T_{1},T_{2},\ldots,T_{k}\subseteq\Omega$ be sets with membership oracles $O_{1},\ldots,O_{k}$ respectively. Let $\epsilon_{\textsf{sz}}\geq 0$ and let $\textsf{sz}_{1},\ldots,\textsf{sz}_{k}\in\mathbb{N}$ be such that for every $i\leq k$, we have $\frac{\left|T_{i}\right|}{(1+\epsilon_{\textsf{sz}})}\leq\textsf{sz}_{i}\leq(1+\epsilon_{\textsf{sz}})\left|T_{i}\right|$. For each $i\leq k$, let $S_{i}$ be a sequence of samples (with possible repetitions) obtained by repeatedly sampling elements from $T_{i}$ uniformly with replacement such that $|S_{i}|\geq\texttt{thresh}$, where $\texttt{thresh}=24\cdot\frac{(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\cdot\log(\frac{4k}{\delta})$ Then, on input $\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$, the algorithm AppUnion with parameters $\epsilon,\delta$ (Algorithm 2) outputs sz that satisfies:
${\mathsf{Pr}}\!\left[\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}\leq\left|\textsf{sz}\right|<(1+\epsilon)(1+\epsilon_{\textsf{sz}})|\cup_{i=1}^{k}T_{i}|\right]>1-\delta.$
The worst case number of calls to the oracles $O_{i}$ made by the algorithm is $O\left(k\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot\frac{1}{\epsilon^{2}}\cdot\log\big{(}\frac{k}{\delta}\big{)}\right)$ and the rest of the computation takes a similar amount of time.
[…]
We next tread towards formally characterizing the guarantee of Algorithm 3. To this end, we will introduce few notations:
  1. (1)
    For each $1\leq i\leq\ell$ and each $q\in Q$, let $\texttt{AccurateN}_{i,q}$ be the event that $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|$. Also, let $\texttt{AccurateN}_{i}=\bigcap\limits_{q\in Q}\texttt{AccurateN}_{i,q}$ and $\texttt{AccurateN}_{\leq i}=\bigcap\limits_{1\leq j\leq i}\texttt{AccurateN}_{j}$.
  2. (2)
    For a state ${q}^{\ell}$, we will use $\mathsf{U}({q}^{\ell})$ to denote the random variable denoting the sequence of samples (with possible repetitions) constructed by repeatedly sampling $|\mathsf{S}({q}^{\ell})|$ number of elements uniformly from $\mathcal{L}({q}^{\ell})$. Our algorithms compute sequences $\mathsf{S}({q}^{\ell})$ of elements (strings) sampled from $\mathcal{L}({q}^{\ell})$. For arguing about correctness of our algorithms, we will in turn argue that the distribution of $\mathsf{S}({q}^{\ell})$ would be “similar" to that of $\mathsf{U}({q}^{\ell})$. Towards this, we will often invoke the the probability space obtained by replacing the sets $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$. Formally, for any event $\mathcal{E},\mathcal{F}$ (in the context of Algorithm 3 and Algorithm 4) we will denote by ${\mathsf{Pr}}\!\left[\mathcal{E};{{\mathsf{U}}_{\leq i}}\right]$ (resp. ${\mathsf{Pr}}\!\left[\mathcal{E}\;|\;\mathcal{F};{{\mathsf{U}}_{\leq i}}\right]$) the probability of the event $\mathcal{E}$ (resp. probability of the event $\mathcal{E}$, given event $\mathcal{F}$) when the $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 is replaced by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$.
  3. (3)
    Consider the random trial corresponding to the call to the procedure $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$ for $\ell>0$ and let $\mathtt{ret}_{\textnormal{{sample}}}$ denote the random variable representing the return value thus obtained. Observe that, in each such trial, the function AppUnion is called $2\ell$ times, twice for each level $1\leq i\leq\ell$. Let us now define some events of interest.
  4. (4)
    For each $1\leq i\leq\ell$ and ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$, let us denote by $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ the event that the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}^{\text{th}}$ call to AppUnion corresponding to level $i$ returns a value $\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ that satisfies 1($\frac{1}{(1+\beta)^{i}}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|\leq\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq(1+\beta)^{i}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|$. Here, $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$. For ease of notation, we also define $\texttt{Accurate}\texttt{AppUnion}=\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$
  5. (5)
    Let $\mathtt{Fail}_{1}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned because $\varphi>1$ at the time of return (Line 3). Let $\mathtt{Fail}_{2}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned at Line 3. Finally, let $\mathtt{Fail}=\mathtt{Fail}_{1}\uplus\mathtt{Fail}_{2}$.
The following presents the correctness of Algorithm 3, while its proof is presented in Section 5.
Theorem 2.
Let $\gamma_{0},\beta$ and $\{\mathsf{S}({p}^{i})\}_{1\leq i<\ell,p\in Q}$ satisfy the following:
$\begin{array}[]{c}\gamma_{0}=\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})},\quad\quad\beta\leq\frac{1}{4\cdot n^{2}},\quad\quad\text{ and,}\\ \forall\;1\leq i\leq\ell,\forall p\in Q,\;\;|\mathsf{S}({p}^{i})|\geq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\end{array}$
Then for the trial corresponding to the invocation $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$, the following statements hold true
  1. (1)
    For each $w\in\mathcal{L}({q}^{\ell})$,
    ${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\;\;\big{|}\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=\gamma_{0}$
  2. (2)
    ${\mathsf{Pr}}\!\left[\mathtt{Fail}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]\leq 1-\frac{2}{3e^{2}}$
  3. (3)
    ${\mathsf{Pr}}\!\left[\texttt{Accurate}\texttt{AppUnion}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\frac{\eta}{2}$
Input : NFA $\mathcal{A}$ with $m$ states, $n\in\mathbb{N}$ presented in unary
Parameters : Accuracy parameter $\varepsilon$ and confidence parameter $\delta$
1 $\beta\leftarrow\frac{\varepsilon}{4n^{2}}$ $\eta\leftarrow\frac{\delta}{2\cdot n\cdot m}$ $\mathtt{ns}\leftarrow\frac{4096\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{4096\cdot m^{2}\cdot n^{2}\cdot\log(\varepsilon^{-2})}{\delta}\right)$ $\mathtt{xns}\leftarrow\mathtt{ns}\cdot 12\cdot(1-\frac{2}{3e^{2}})^{-1}\cdot\log(8/\eta)$ Construct the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$ from $\mathcal{A}$ input $n$ for $\ell\in\{0,\ldots,n\}\text{ and }q\in Q$ do
2     if $\ell=0$ then
3          $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow(q\in F)$ $\big{(}1,\{\lambda\}\big{)}$ : $\big{(}0,\varnothing)$
4     foreach ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$ do
5          $\{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1},\ldots,p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}\}\leftarrow\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ $\beta^{\prime}=(1+\beta)^{\ell-1}-1$ $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\texttt{AppUnion}_{\beta,\eta/2}\bigg{(}\beta^{\prime},$ $(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1})),\ldots,(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1})),$ $\mathcal{A}_{\textsf{unroll}}\bigg{)}$   // The last parameter $\mathcal{A}_{\textsf{unroll}}$ is a proxy for the membership oracles required by AppUnion
6
7    $\mathsf{N}({q}^{\ell})\leftarrow\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ $\mathsf{S}({q}^{\ell})\leftarrow\varnothing$ repeat $\mathtt{xns}$ times:
8          if $|\mathsf{S}({q}^{\ell})|<\mathtt{ns}$ then
9               $w\leftarrow$ sample(,$\ell,\{{q}^{\ell}\},\lambda,\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})},\beta,\frac{\eta}{2\cdot\mathtt{xns}}$) if $w\neq\bot$ then
10                    $\mathsf{S}({q}^{\ell})\leftarrow\mathsf{S}({q}^{\ell})\boldsymbol{\!\cdot}\!\mathsf{append}(w)$
11
12          else break
13    let $\alpha_{{q}^{\ell}}=\mathtt{ns}{-}|\mathsf{S}({q}^{\ell})|$ if $\alpha_{{q}^{\ell}}>0$ then
14          let $w_{{q}^{\ell}}$ be some word in $\mathcal{L}({q}^{\ell})$ repeat $\alpha_{{q}^{\ell}}$ times:  $\mathsf{S}({q}^{\ell}){\leftarrow}\mathsf{S}({q}^{\ell})\boldsymbol{\!\cdot}\!\mathsf{append}(w_{{q}^{\ell}})$
15
return $\mathsf{N}({q}^{n}_{F})$
Algorithm 4 Algorithm for Estimating $|\mathcal{L}(\mathcal{A})_{n}|$
The main algorithm. Algorithm 4 first constructs (in Line 4) the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$. The algorithm then goes over all the states and all levels $0\leq\ell\leq n$ to construct the sampled sets and the size estimates inductively. In Lines 4-4, the algorithm caters for the base case ($\ell=0$). In the inductive case (Lines 4-4) the algorithm computes the estimate $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$, where $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the estimate of the size of $\cup_{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}\in\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}}^{\ell-1})$. The size estimates of, and a sample subsets of the sets $\{\mathcal{L}({q}^{\ell-1})\}_{q\in Q}$ are available inductively at this point. Further, membership in these languages can be easily checked. As a result, the algorithm uses the subroutine AppUnion to compute the size of the union (in Line 4). Once the size estimates $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ are determined, the algorithm constructs the sampled subset of $\mathcal{L}({q}^{\ell})$ by making $\mathsf{xns}$ calls to the subroutine sample (Lines 4-4). Theorem 3 presents the desired correctness statement of our FPRAS  Algorithm 4, and its proof is presented in Section 6.
Theorem 3.
Given an NFA $\mathcal{A}$ with $m$ states and $n\in\mathbb{N}$ (presented in unary), Algorithm 4 returns $\mathsf{Est}$ such that the following holds:
${\mathsf{Pr}}\!\left[\frac{|L(F^{n})|}{1+\epsilon}\leq\mathsf{Est}\leq(1+\epsilon)|L(F^{n})|\right]\geq 1-\delta$
Moreover, the algorithm has time complexity $\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$, where the tilde hides for polynomial factors of $\log(m+n)$.
4. Correctness of Algorithm 2
The proof is in the same lines as that of Karp and Luby (Karp and Luby, 1985); for completeness, we present a full proof.
Proof.
Let $\mathsf{Fail}$ be the event that the output of the algorithm is not within $\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$ and $|\cup_{i=1}^{k}T_{i}|{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$. Clearly, we need to upper bound $\Pr[\mathsf{Fail}]$ by $\delta$.
Note that the algorithm in each of the $t$ runs of the loop (lines 2-2) tries to draw an element from some $S_{i}$. So if we assume that the size of the sets $S_{i}$ for all $i$ is more than $t$, then the condition is the if (in line 2 will be satisfied and in that case the else clause in line 2 to 2 is redundant.
We will prove the correctness of  Algorithm 2 in two parts. In the first part we will prove that if the size of the sets $S_{i}$ is greater than $t$ then $\Pr[\mathsf{Fail}]\leq\frac{\delta}{2}$. In the second part we will prove that if the sets $S_{i}$ has size thresh (much smaller than $t$) then the probability that the algorithm will ever need the else clause in line 2 to 2 is less that $\delta/2$. The two parts combined proves that if the sets $S_{i}$s are of size threshthen $\Pr[\mathsf{Fail}]\leq\delta$ which is what is claimed in the theorem.
For all $\sigma\in\cup_{i=1}^{k}T_{i}$ let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$. In other words, $\sigma$ is in $T_{L(\sigma)}$ and for all $i<L(\sigma)$, $\sigma\not\in T_{i}$. Observe $\mathcal{U}_{1}:=\left\{(\sigma,L(\sigma)\ \mid\ \sigma\in\cup_{i=1}^{k}T_{i}\right\}$ has size exactly $\left|\cup_{i=1}^{k}T_{i}\right|$. Furthermore, $\mathcal{U}_{2}:=\left\{(\sigma,i)\ \mid\ \sigma\in T_{i}\right\}$ has size $\sum_{i=1}^{k}|T_{i}|$.
Part 1: If $\forall i$, $|S_{i}|\geq t$ then $\Pr[\mathsf{Fail}]\leq\frac{\delta}{2}$.
Note that in this case the condition in the if clause in line 2 is always satisfied. For a run of the loop (lines 2-2), consider the lines 2 to 2. We say the pair $(\sigma,i)$ is sampled if in one of the iteration of the loop $i$ is sampled in line 2 and then in that same iteration $\sigma$ is obtained from $S_{i}.\mathsf{deque()}$ in the line 2.
For any $i_{0}\in[t]$ and $\sigma_{0}\in S_{i_{0}}$, we have
[…]
$=$ $\frac{\textsf{sz}_{i_{0}}}{\sum_{i}^{k}\textsf{sz}_{i}}\times\frac{1}{|T_{i_{0}}|}$
The last equality is because we have assumed $\forall i$, $|S_{i}|\geq t$, so the if condition in line 2 is always satisfied. Hence once $i_{0}$ is sampled in line 2 probability that $\sigma_{0}$ is obtained is exactly $1/|T_{i_{0}}|$, since the set $S_{i_{0}}$ contains elements sampled uniformly (with replacement) from $T_{i_{0}}$. Therefore,
$\frac{1}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\Pr[(\sigma_{0},i_{0})\mbox{ is sampled }]\leq(1+\epsilon_{\textsf{sz}})\frac{1}{\sum_{i=1}^{k}\textsf{sz}_{i}}.$
Let $Y_{j}$ be the random variable denoting whether the counter $Y$ is increased in the $j$th iteration of the loop. Note that $Y_{j}=1$ if the pair $(\sigma_{0},i_{0})$ sampled in the $j$th iteration is in $\mathcal{U}_{1}$. This is because in line 2 it is checked if the sampled pair is in $\mathcal{U}_{1}$.
So, $\mathbb{E}[Y_{j}]=\sum_{(\sigma,i)\in\mathcal{U}_{1}}\Pr[(\sigma,i)\mbox{ is sampled}]$. Thus,
$\frac{|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\mathbb{E}[Y_{j}]\leq(1+\epsilon_{\textsf{sz}})\frac{|\mathcal{U}_{1}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$
Note that $Y=\sum_{i}^{t}Y_{i}$. Thus at the end of the algorithm the expected value of $Y$ is between $\frac{t|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}$ and $(1+\epsilon_{\textsf{sz}})^{2}\frac{t|\mathcal{U}_{1}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$.
Now by Chernoff bound we have
$\Pr[Y\not\in(1\pm\frac{\epsilon}{2})\mathbb{E}[Y]]\leq 2exp(-\frac{\epsilon^{2}\mathbb{E}[Y]}{12}).$
We also know that,
$\mathbb{E}[Y]\geq$ $\frac{t|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}|T_{i}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}\textsf{sz}_{i}}{(1+\epsilon_{\textsf{sz}})^{2}\sum_{i=1}^{k}\textsf{sz}_{i}}$
$\geq$ $12\log(\frac{4}{\delta})$
The last inequality is due to the setting of the parameter $t$. Thus, by Chernoff bound, probability that the output of the algorithm is not between $(1-\frac{\epsilon}{2})\frac{|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})}$ and $(1+\frac{\epsilon}{2})(1+\epsilon_{\textsf{sz}})|\mathcal{U}_{1}|$ is at most $\delta/2$. Since $(1-\frac{\epsilon}{2})\geq\frac{1}{1+\epsilon}$, so we have $\Pr[\mathsf{Fail}\ \cap\ \overline{\mathsf{Bad}}]\leq\frac{\delta}{2}$.
Part 2: If for all $i$, $|S_{i}|\geq\texttt{thresh}$
Note that, the else clause in line 2 to 2 is ever reached, if for some $i$, the number of times $i$ is sampled in the $t$ iterations is more than thresh.
Let $\mathsf{Bad}_{i}$ denote the event that the index $i$ is sampled (in line 2) more that thresh number of times during the course of the algorithm, that is during the $t$ runs of the loop (lines 2-2). And let the event $\mathsf{Bad}$ be $\cup_{i=1}^{k}\mathsf{Bad}_{i}$. We will now upper bound, $\Pr[\mathsf{Bad}_{i}]$. Let $W_{i}$ be the random variable counting the number times $i$ is sampled. $\mathbb{E}[W_{i}]=t\times\frac{\textsf{sz}_{i}}{\sum_{i=1}^{k}\textsf{sz}_{i}}$ which is less than $\frac{12(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\log(4/\delta)$. By simple application of Chernoff Bound we see that
$\Pr[W_{i}\geq(1+\Delta)\mathbb{E}[W_{i}]]\leq exp(-\frac{\Delta^{2}\mathbb{E}[W_{i}]}{2+\Delta})\leq exp(-\frac{\Delta\mathbb{E}[W_{i}]}{2}),$
the last inequality is for $\Delta\geq 1$. For our purpose if we put $\Delta=\log(\frac{2k}{\delta})\frac{1}{\mathbb{E}[W_{i}]}$, since $|S_{i}|\geq\texttt{thresh}\geq\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ + $\log(\frac{2k}{\delta})$, we have $\Pr[\mathsf{Bad_{i}}]\leq\frac{\delta}{2k}$. So by union bound $\Pr[\mathsf{Bad}]\leq\delta/2$.
5. Correctness of Algorithm 3
In the following three sections, we prove the three parts of Theorem 2 individually.
5.1. Proof of Theorem 2(1)
Consider an execution of Algorithm 3 and let $w=w_{1}w_{2}\ldots w_{\ell}\in\mathcal{L}({q}^{\ell})$ be the string constructed right before entering the branch at Line 3. Let $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}$ be the value of the variable $\mathsf{pr_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}$ at level $i$ and let $v$ be the value of $\varphi$ before the function returns. Then we have,
$v=\frac{\gamma_{0}}{p_{w_{\ell},\ell}\cdot p_{w_{\ell-1},\ell-1}\cdots p_{w_{1},1}}$
Let $\textsf{sz}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},\textsf{sz}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},\textsf{sz}^{\ell-1}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},\textsf{sz}^{\ell-1}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},\ldots\textsf{sz}^{1}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},\textsf{sz}^{1}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ be the estimates obtained in the $2\ell$ calls to AppUnion during the run of $\textnormal{{sample}}()$. Then,
$p_{w_{i},i}=\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}$
Thus,
(2) $v=\frac{\gamma_{0}}{\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}}=\gamma_{0}\cdot\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}{\textsf{sz}^{i}_{w_{i}}}$
Recall that $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$. Observe that $\mathcal{L}(P^{i})=\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})\uplus\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})$, and thus,
$|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})|+|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})|=|\mathcal{L}(P^{i})|$
Hence, under the event $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}\cap\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ we have
(3) $\frac{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}{\textsf{sz}^{i}_{w_{i}}}\leq(1+\beta)^{2i}\cdot\frac{|\mathcal{L}(P^{i})|}{|\mathcal{L}(P^{i-1})|}$
This gives us,
$\begin{array}[]{c}v\leq\gamma_{0}\cdot\prod\limits_{i=1}^{\ell}(1+\beta)^{2i}\cdot\frac{|\mathcal{L}(P^{i})|}{|\mathcal{L}(P^{i-1})|}=\gamma_{0}{\cdot}\Bigg{(}\prod\limits_{i=1}^{\ell}(1+\beta)^{2i}\Bigg{)}{\cdot}\Bigg{(}\prod\limits_{i=1}^{\ell}\frac{|\mathcal{L}(P^{i})|}{|\mathcal{L}(P^{i-1})|}\Bigg{)}\end{array}$
Now, observe that $|\mathcal{L}(P^{0})|=1$ and $P^{\ell}=\{{q}^{\ell}\}$. Thus, under the event $\big{(}\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\big{)}=\texttt{Accurate}\texttt{AppUnion}$, we have
(5) $\begin{array}[]{c}v\leq\gamma_{0}{\cdot}\Bigg{(}\prod\limits_{i=1}^{\ell}(1+\beta)^{2i}\Bigg{)}{\cdot}|\mathcal{L}({q}^{\ell})|=\gamma_{0}{\cdot}(1+\beta)^{\ell(\ell+1)}{\cdot}|\mathcal{L}({q}^{\ell})|\end{array}$
Assuming the event $\texttt{AccurateN}_{\ell,q}$ also holds, we have $\frac{|\mathcal{L}({q}^{\ell})|}{\mathsf{N}({q}^{\ell})}\leq(1+\beta)^{\ell}$. Under this assumption, substituting $\gamma_{0}=\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})}$, $\beta\leq\frac{1}{4\cdot n^{2}}$, and using the inequality $(1+1/4n^{2})^{n}\leq e$, we have,
$\begin{array}[]{c}v\leq\frac{2}{3e}{\cdot}(1+\beta)^{\ell(\ell+2)}\leq\frac{2}{3e}\cdot e\leq 1\\ \end{array}$
It then follows that
${\mathsf{Pr}}\!\left[\mathtt{Fail}_{1}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=0$
Substituting, $v=\frac{\gamma_{0}}{p_{w_{\ell},\ell}\cdot p_{w_{\ell-1},\ell-1}\cdots p_{w_{1},1}}$, we then have
$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]\\ =v\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\frac{\gamma_{0}}{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},\ell}\cdot p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},\ell-1}\cdots p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},1}}\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\gamma_{0}\end{array}$
5.2. Proof of Theorem 2(2)
We will now estimate the (conditional) probability of the event $\mathtt{Fail}$. First, observe that for each string $w^{\prime}\notin\mathcal{L}({q}^{\ell})$, we have ${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w^{\prime}\right]=0$ because any string $w$ that the algorithm outputs is such that there is a path from the initial state to ${q}^{\ell}$ labeled with $w$, and (b) the unrolled automata is such that any path from start state to some state ${p}^{i}$ corresponds to a string in $\mathcal{L}({p}^{i})$.
We observe that $\overline{\mathtt{Fail}}$ is the event that $\mathtt{ret}_{\textnormal{{sample}}}$ is a string in $\mathcal{L}({q}^{\ell})$, each of which is a disjoint event. Thus,
${\mathsf{Pr}}\!\left[\mathtt{Fail}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]$
$=1{-}\sum\limits_{w\in\mathcal{L}({q}^{\ell})}{\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]$
$=1{-}\sum\limits_{w\in\mathcal{L}({q}^{\ell})}\gamma_{0}=1{-}\gamma_{0}|\mathcal{L}({q}^{\ell})|\leq 1-\frac{2}{3e}\cdot\frac{|\mathcal{L}({q}^{\ell})|}{\mathsf{N}({q}^{\ell})}$
$\leq 1-\frac{2}{3e}\cdot\frac{1}{(1+\beta)^{\ell}}\leq 1-\frac{2}{3e^{2}}$
5.3. Proof of Theorem 2(3)
We would like to use Theorem 1 to prove Theorem 2(3). Towards this, we first establish that the pre-conditions of Theorem 1 hold.
  1. (1)
    For $i\leq\ell$, if we substitute $\beta^{\prime}=(1+\beta)^{i-1}-1,\eta^{\prime}=\eta/{4n}$, we get $\texttt{thresh}=24\cdot\frac{(1+\beta)^{2i-2}}{\beta^{2}}\cdot\log\left(\frac{16k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}n}{\eta}\right)$, where $k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the number of sets in the argument of AppUnion. Observe that, $k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq m$ and that, for each $i\leq\ell$, we have $(1+\beta)^{2i-2}\leq(1+\beta)^{4n^{2}}\leq e$. Thus for each ${q}^{i}$, we have
    $\texttt{thresh}\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\leq|\mathsf{S}({q}^{i-1})|$
  2. (2)
    Since we are considering the distribution obtained when $\mathsf{S}({q}^{i})$ is replaced with $\mathsf{U}({q}^{i})$ for each $i\leq\ell-1$ (observe the ${\mathsf{Pr}}\!\left[;{{\mathsf{U}}_{\leq\ell-1}}\right]$ in the statement of Theorem 2(3)), we directly have that each input $\{S_{i}\}_{i}$ to every call to AppUnion has at least thresh elements (as established) elements obtained by repeatedly sampling $\{T_{i}\}_{i}$ uniformly.
  3. (3)
    The event $\texttt{AccurateN}_{\leq\ell-1}$ demands that for each $i\leq\ell-1$, we have $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|$.
From Theorem 1, it thus follows that, for all $i$,
${\mathsf{Pr}}\!\left[\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta^{\prime}$
Therefore, applying union-bound, we have
$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\\ \geq 1-2n\cdot\eta^{\prime}=1-\frac{\eta}{2}\end{array}$
6. Correctness of Algorithm 4
We first establish some preliminaries relevant to our proof.
6.1. Total Variation Distance
The total variation distance is a statistical distance between two discrete valued random variables $X,Y$ over domain $\Omega$ and is defined as follows:
$\mathsf{TV}(X,Y)$ $=$ $\frac{1}{2}\sum_{v\in\Omega}\left|\Pr[X=v]-\Pr[Y=v]\right|$
$=$ $\sum_{v\,:\,\Pr[X=v]\geq\Pr[Y=v]}\left(\Pr[X=v]-\Pr[Y=v]\right)\;.$
We will use the following folklore result (c.f. Lemma 6.3 of  (Vadhan, 2012)). Here, for a random variable $X$ over the sample space $\Omega$, we use the notation ${\mathsf{Pr}}\!\left[E:X\right]$ to denote the probability of the event $E\subseteq\Omega$.
Lemma 6.0.
Let $X$ and $Y$ be the random variables taking values over a measurable set $\Omega$. Let $f:\Omega\to\hat{\Omega}$ be a (possibly randomized) function with measurable co-domain $\hat{\Omega}$. Let $E\subseteq\hat{\Omega}$ be an arbitrary event. Then, the following holds:
$|{\mathsf{Pr}}\!\left[E:{f(X)}\right]-{\mathsf{Pr}}\!\left[E:f(Y)\right]|\leq\mathsf{TV}(X,Y)$
6.2. Proof of Theorem 3
We prove the main result (Theorem 3) by induction on the level $\ell$. For each level $\ell$, we first characterize the accuracy of the computed estimates $\{\mathsf{N}({q}^{\ell})\}_{q\in Q}$ and the quality of the sampled sets $\{\mathsf{S}({q}^{\ell})\}_{q\in Q}$, assuming that these were computed when the samples at the lower levels $11,2,\ldots,\ell-$ 1 are perfectly uniform and have size thresh. After this, we characterize the real computed estimates and samples by arguing that the distance of the corresponding random variables, from those computed using perfectly uniform samples, is small. We will be using the total variation distance and Lemma 1 to structure the proof. We first establish some helpful auxiliary results (Lemma 2 and Lemma 3) towards our main proof. As before, we will often invoke the the probability space obtained by replacing the sets $\mathsf{S}({q}^{j})$ by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$. In addition to the previously introduced notation ${\mathsf{Pr}}\!\left[;{{\mathsf{U}}_{\leq i}}\right]$, we will also use the notation ${\mathsf{TV}}(X,Y\;;{\mathsf{U}}_{\leq i})$ (resp. ${\mathsf{TV}}(X,Y\;\;\big{|}\;\;\mathcal{F}\;;{\mathsf{U}}_{\leq i})$) to denote the total variational distance between the distribution induced by random variables $X,Y$ (resp. distribution induced by $X,Y$ conditioned on the event $\mathcal{F}$) when the $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 is replaced by a set $\mathsf{U}({q}^{j})$ obtained by unirofmly sampling (with replacement) $|\mathsf{S}({q}^{j})|$ many elements from $\mathcal{L}({q}^{j})$, for all $q\in Q$ and for all $1\leq j\leq i$.
Lemma 6.0.
For each $\ell\leq n$ and $q\in Q$, we have
${\mathsf{Pr}}\!\left[\texttt{AccurateN}_{q,\ell}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta$
Proof.
[…]
thresh $=24\cdot\frac{(1+\beta)^{2\ell-2}}{\beta^{2}}\cdot\log\left(\frac{4k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}{\eta^{\prime}}\right)\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{8m^{2}n}{\delta}\right)$
$\leq\frac{384\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{8m^{2}}{\delta}\right)\leq\mathtt{ns}$
Therefore, from Theorem 1, we have
${\mathsf{Pr}}\!\left[\begin{array}[]{l}\frac{|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|}{(1+\beta)^{\ell}}\leq\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\\ \quad\quad\leq(1+\beta)^{\ell}\;|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|\end{array}\big{|}\;\;\texttt{AccurateN}_{\ell-1};{{\mathsf{U}}_{\ell-1}}\right]\geq 1-\frac{\eta}{2}$
Since $\mathsf{N}({q}^{\ell})=\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$, we have
${\mathsf{Pr}}\!\left[\begin{array}[]{l}\frac{|\mathcal{L}({q}^{\ell})|}{(1+\beta)^{\ell}}\leq\mathsf{N}({q}^{\ell})\\ \quad\quad\leq(1+\beta)^{\ell}|\mathcal{L}({q}^{\ell})|\end{array}\mid\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta$
Lemma 6.0.
For each $\ell\leq n$ and $q\in Q$, we have
${\mathsf{TV}}(S(q^{\ell}),U(q^{\ell})\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\ell-1})\leq\eta$
Proof.
Fix $\ell$ and $q$. Let $\mathtt{Fail}^{j}$ denote the event that the $j^{\text{th}}$ call to sample (in the loop starting at Line 4) for $i=j$ returns $\bot$. Similarly, let $\texttt{Accurate}\texttt{AppUnion}^{j}$ denote the event AccurateAppUnion corresponding to the call to sample for $i=j$. Using Iversion bracket notation, we denote the indicator variable $[\mathtt{Fail}^{j}]$. Also, we use SmallS to denote the event that $\alpha_{{q}^{\ell}}>0$. Observe that SmallS occurs if and only if $\sum\limits_{j=1}^{\mathtt{xns}}[\mathtt{Fail}^{j}]\geq\mathtt{xns}-\mathtt{ns}$.
Recall, we have, for each $j$
${\mathsf{Pr}}\!\left[[\mathtt{Fail}^{j}]\;\;\big{|}\;\;\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq 1-\frac{2}{3e^{2}}$
Therefore, from Chernoff bound, we have
${\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\;\;\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\frac{\eta}{2}$
Furthermore, from Theorem 2 (part 3), we have
${\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\eta/2$
Note that assuming $\texttt{AccurateN}_{\leq\ell}$ and ${\mathsf{U}}_{\leq\ell-1}$, is the event SmallS does not happen and for all $j$ $\texttt{Accurate}\texttt{AppUnion}^{j}$ holds then the distribution of $S(q^{\ell})$ and $U(q^{\ell})$ is identical. So,
$\mathsf{TV}(S(q^{\ell}),U(q^{\ell})\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell}\;;\;{\mathsf{U}}_{\leq\ell-1})$
$\leq{\mathsf{Pr}}\!\left[\textsc{SmallS}\cup\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\leq{\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\quad+{\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\leq\eta/2+\eta/2=\eta$
Now, we are ready to state and prove the main theorem.
Proof of Theorem 3.
For ease of exposition, we will refer to the event $\texttt{AccurateN}_{n}$ in our analysis. Applying  Lemma 1, we have
${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq$
$\sum_{\ell\in[1,n];q\in Q}{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\qquad+\sum\limits_{\ell\in[1,n-1];q\in Q}\mathsf{TV}(S(q^{\ell}),U(q^{\ell})\mid\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
By Lemma 2 and Lemma 3 and applying union bound oven all $\ell\in[1,n]$ and all $q\in Q$ we have
${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq n\cdot m\cdot\eta+n\cdot m\cdot\eta\leq\delta.$
To calculate the time complexity of Algorithm 4 we first observe that the for loop goes over $nm$ elements. Now for each $\ell$ and $q$ there is one call to $\texttt{AppUnion}_{\beta,\eta}$ with the number of sets being $O(m)$ and $\mathtt{xns}$ calls to sample(). So by Theorem 1 the time taken by AppUnion is $O(m/\beta^{2}\log(m/\eta))$ number of calls to the oracle. So the total complexity is $\widetilde{O}(mn^{4}/\epsilon^{2}\log(1/\delta))$, ignoring the factors of $\log(n+m)$. The sample() on the other hand is a recursive function which calls $\texttt{AppUnion}_{\beta,\eta}()$ every recursion and the recursion happens $O(n)$ times. So the time complexity for each of the $\mathtt{xns}$ calls to sample() is $\widetilde{O}(mn^{5}/\epsilon^{2}\log(1/\delta))$ calls to the oracle. Now note that the membership oracle takes $O(m^{2}n)$ steps for each call. So in total the time complexity of the Algorithm 4 is $\widetilde{O}(mn\cdot\mathtt{xns}\cdot\frac{mn^{5}}{\epsilon^{2}}\log(1/\delta)\cdot(m^{2}n))=\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.
7. Conclusions and Future Work
We consider the approximate counting problem #NFA and propose a fully polynomial time randomized approximation scheme (FPRAS) that significantly improves the prior FPRAS (Arenas et al., 2019). Our algorithm deeply exploits recent progress in the streaming algorithms literature. Given the wide range of applications of counting and sampling from the language of an NFA in program analysis, testing and more broadly in Computer Science, we envision that further improvements in the complexity of approximating #NFA is a worthwhile avenue for future work.
Acknowledgements.
The authors thank Pierre Snellart and Antoine Amarilli for introducing us to the #NFA problem, and are grateful to N.V. Vinodchandran and Pavan Aduri for insightful discussions on the Karp-Luby method for the union of sets. This work was supported in part by National Research Foundation Singapore under its NRF Fellowship Programme[NRF-NRFFAI1-2019-0004 ], Ministry of Education Singapore Tier 2 grant MOE-T2EP20121-0011, and Ministry of Education Singapore Tier 1 grants [R-252-000-B59-114] and [R-252-000-C28-133].

### 08 Apr 2024 v2

@@
 While the corresponding decision problem of checking the emptiness of $\mathcal{L}(\mathcal{A}_{n})$ is solvable in polynomial time, the #NFA problem is known to be #P-hard.
-Recently, the long-standing open question — whether there is an FPRAS (fully polynomial time randomized approximation scheme) for #NFA — was resolved in (Arenas et al., 2019).
-The FPRAS due to (Arenas et al., 2019) relies on the interreducibility of counting and sampling, and computes, for each pair of state $q$ and natural number $i\leq n$, a set of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ many uniformly chosen samples from the set of words of length $i$ that have a run ending at $q$.
-This informative measure — the number of samples maintained per state and length — also affects the overall time complexity with a quadratic dependence.
-Given the prohibitively high time complexity, in terms of each of the input parameters, of the FPRAS due to (Arenas et al., 2019), and considering the widespread application of approximate counting (and sampling) in various tasks in Computer Science, a natural question arises: Is there a faster FPRAS for #NFA that can pave the way for the practical implementation of approximate #NFA tools?
-In this work, we demonstrate that significant improvements in time complexity are achievable.
-Specifically, we have reduced the number of samples required for each state to be independent of $m$, with significantly less dependence on $n$ and $\epsilon$, maintaining only $\widetilde{O}(\frac{n^{4}}{\epsilon^{2}})$ samples per state.
-We believe our insights may lead to further reductions in time complexity and thus open up a promising avenue for future work towards the practical implementation of tools for approximating #NFA.
-The authors decided to forgo the old convention of alphabetical ordering of authors in favor of a randomized ordering, denoted by r⃝.
-The publicly verifiable record of the randomization is available at https://www.aeaweb.org/journals/policies/random-author-order/search
-copyright: noneconference: ; ;
+Recently, the long-standing open question — whether there is an FPRAS (fully polynomial time randomized approximation scheme) for #NFA — was resolved by Arenas, Croquevielle, Jayaram, and Riveros in (Arenas et al., 2019).
+The authors demonstrated the existence of a fully polynomial randomized approximation scheme with a time complexity of $\widetilde{O}\left(m^{17}n^{17}\cdot\frac{1}{\epsilon^{14}}\cdot\log(1/\delta)\right)$, for a given tolerance $\epsilon$ and confidence parameter $\delta$.
+Given the prohibitively high time complexity in terms of each of the input parameters, and considering the widespread application of approximate counting (and sampling) in various tasks in Computer Science, a natural question arises: is there a faster FPRAS for #NFA that can pave the way for the practical implementation of approximate #NFA tools?
+In this work, we answer this question in the positive.
+We demonstrate that significant improvements in time complexity are achievable, and propose an FPRAS for #NFA that is more efficient in terms of both time and sample complexity.
+A key ingredient in the FPRAS due to Arenas, Croquevielle, Jayaram, and Riveros (Arenas et al., 2019) is inter-reducibility of sampling and counting, which necessitates a closer look at the more informative measure — the number of samples maintained for each pair of state $q$ and length $i\leq n$.
+In particular, the scheme of (Arenas et al., 2019) maintains $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ samples per pair of state and length.
+In the FPRAS we propose, we systematically reduce the number of samples required for each state to be only poly-logarithmically dependent on $m$, with significantly less dependence on $n$ and $\epsilon$, maintaining only $\widetilde{O}(\frac{n^{4}}{\epsilon^{2}})$ samples per state.
+Consequently, our FPRAS runs in time $\widetilde{O}((m^{2}n^{10}+m^{3}n^{6})\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.
+The FPRAS and its analysis use several novel insights.
+First, our FPRAS maintains a weaker invariant about the quality of the estimate of the number of samples for each state $q$ and length $i\leq n$.
+Second, our FPRAS only requires that the distribution of the samples maintained is close to uniform distribution only in total variation distance (instead of maximum norm).
+We believe our insights may lead to further reductions in time complexity and thus open up a promising avenue for future work towards the practical implementation of tools for approximate #NFA.
+##\##NFA, counting, estimation, FPRAS
+copyright: rightsretainedjournal: PACMMODjournalyear: 2024journalvolume: 2journalnumber: 2 (PODS)article: 112publicationmonth: 5doi: 10.1145/3651613ccs: Theory of computation Theory and algorithms for application domains
 1. Introduction
@@
 : Given an NFA $\mathcal{A}=(Q,I,\Delta,F)$ over the binary alphabet $\Sigma=\{0,1\}$ with $m$ states, and a number $n$ in unary, determine $|\mathcal{L}(\mathcal{A}_{n})|$, where $\mathcal{L}(\mathcal{A}_{n})$ is the set of strings of length $n$ that are accepted by $\mathcal{A}$.
-The counting problem #NFA and the closely related problem of uniform sampling words of an NFA are fundamental to database applications, and arise when counting and uniformly generating solutions to queries.
-Beyond database systems, the aforementioned problem finds applications in program verification (Legay et al., 2010), testing of software (Sutton et al., 2007) and distributed systems (Ozkan et al., 2019), probabilistic programming, estimation of information leakage in software (Bang et al., 2016; Gao et al., 2019; Saha et al., 2023), and even music generation (Donzé et al., 2014).
-How do we count the number of elements of $\mathcal{L}(\mathcal{A}_{n})$?
-When $\mathcal{A}$ is a DFA with $m$ states, this question can be solved using a simple dynamic programming algorithm that inductively computes, for every $i\leq n$ and for every state $q$ of $\mathcal{A}$, the number of words of length $i$ that end in state $q$: $\mathsf{count}(q,i)=\sum_{b\in\Sigma}\sum_{p:q=\Delta(p,b)}\mathsf{count}(p,i-1)$ taking overall $O(mn)$ time (assuming $|\Sigma|=O(1)$).
-This simple algorithm however fails to work when $\mathcal{A}$ is a non-deterministic automaton as the same word maybe accepted using many different runs.
-Determinizing $\mathcal{A}$ first and then using the same algorithm as for DFAs would give an overall running time of $O(n\cdot 2^{m})$ which is exponential in the size of the description of the input.
+The problem of #NFA is a fundamental problem in Computer Science, with range of applications in database and information extraction systems; we outline a few below.
+Probabilistic Query Evaluation.
+Given a query $Q$ and a database $D$, the problem of probabilistic query evaluation (PQE) is to determine the probability of the query holding on a randomly sampled database, in which each row is included independently with probability given by its annotated value.
+PQE is known to be #P-hard, even in data complexity setting.
+For a large class of queries, PQE can directly be reduced to the #NFA problem.
+In particular, when the schema of the database $D$ consists only of binary relations, and when the query $Q$ is a self-join-free path query, then the PQE problem for $D$ and $Q$ reduces to an instance of #NFA, whose size is linear in the size of $Q$ and linear in the size of $D$ (van Bremen and Meel, 2023), and further, the time to compute the reduction is also linear in $D$ and $Q$.
+Thus an efficient algorithm for solving #NFA directly yields an efficient algorithm for probabilistic query evaluation.
+Counting Answers to Reglar Path Queries.
+Regular path queries (or property paths(Angles et al., 2017) form a rich fragment of graph query languages such as SPARQL designed for graph database systems.
+Here, one abstracts information on edges in the graph database using an alphabet $\Sigma$ and asks a variety of questions about the set of paths that start from a given source node $u$, end at a target node $v$.
+In particular, a regular path query $(uRv)$ asks to enumerate, count or uniformly sample the set of paths (bounded in length by some fixed number $n$) from $u$ to $v$ whose labels match the regular expression $R$.
+Therefore, the problem of counting the number of answers to the queries reduces to the #NFA problem, for the NFA obtained by the cross product of NFA represented by the database (with $u$ as the intial state and $v$ as the final state) and the NFA that $R$ gets compiled down to.
+This reduced instance is again linear in the size of the database as well as the query, and the reduction takes the same time (Arenas et al., 2019).
+Probabilistic Graph Homomorphism.
+A probabilistic graph is a pair $(H,\pi)$, where $H$ is a graph and $\pi$ is the associated probability labeling over edges.
+The associated probability space is defined as the set of all subgraphs of $H$ wherein every edge $e$ is sampled independently with probability $\pi(e)$.
+Given a query Graph $G$ and a probabilistic graph $(H,\pi)$ the problem of probabilistic graph homomorphism problem asks to to compute the probability that a randomly sampled subgraph of $H$ admits a homomorphism from G.
+In case of 1-way (and 2-way) path queries, the problem was again shown to reduce to #NFA (Amarilli et al., 2024).
+Beyond databases.
+From a complexity theoretic viewpoint, #NFA is SpanL-complete and therefore, finds applications in domains beyond databases such as program verification (Legay et al., 2010), testing of software (Sutton et al., 2007) and distributed systems (Ozkan et al., 2019), probabilistic programming, estimation of information leakage in software (Bang et al., 2016; Gao et al., 2019; Saha et al., 2023), information extraction (Arenas et al., 2021), automated reasoning using BDDs (Arenas et al., 2021), and even music generation (Donzé et al., 2014).
+The direct applications of the #NFA problem brings up the central computational question — how do we count the number of elements of $\mathcal{L}(\mathcal{A}_{n})$?
 It turns out that #NFA is, in fact, provably computationally intractable — it is known to be a #P-hard problem (Álvarez and Jenner, 1993).
-Given the intractability, and given the widespread applications of this problem, the next natural question then arises — can we approximate #NFA using a polynomial time algorithm?
-One of the promising paradigms for deriving efficient algorithms for approximating the answer to an intractable problem is the use of randomization to obtain a fully polynomial time randomized approximation scheme (FPRAS).
-Formally, an FPRAS for a real-valued function $f:\mathcal{D}\to\mathbb{N}$ over some domain $\mathcal{D}$ is a randomized algorithm $A:\mathcal{D}\times(0,1)\times(0,1)\to\mathbb{N}$ that for every input $x\in\mathcal{D}$ and parameters $\epsilon\in(0,1)$ and $\delta\in(0,1)$, runs in time $\textsf{poly}(|x|,\frac{1}{\epsilon},\log(\frac{1}{\delta}))$, and satisfies: ${\mathsf{Pr}}\!\left[\frac{1}{(1+\epsilon)}\cdot f(x)\leq|A(x,\epsilon,\delta)|\leq(1+\epsilon)\cdot f(x)\right]\geq 1-\delta$.
+Given the intractability, and given the widespread applications of the #NFA problem, the next natural question then arises — can we approximate #NFA using a polynomial time algorithm?
 Until very recently, the only approximation algorithms known for #NFA had running times that were quasi-polynomial in the size of the inputs (Gore et al., 1997; Kannan et al., 1995).
-In a recent breakthrough by Arenas, Croquevielle, Jayram and Riveros (Arenas et al., 2019, 2021), #NFA was finally shown to admit an FPRAS.
-Even though the problem was shown to admit an FPRAS, the time complexity of the proposed FPRAS is prohibitively large — it takes time $\widetilde{O}\left(m^{17}n^{17}\cdot\frac{1}{\epsilon^{14}}\cdot\log(1/\delta)\right)$, where the notation $\widetilde{O}$ hides the polylogarithmic multiplicative factors in $m$ and $n$.
-To be applicable in a practical setting like estimating information flow in programs, a polynomial time approximation scheme whose time complexity varies as $O(|P|^{16})$, where $|P|$ is roughly the lines of code, is unlikely to scale even for moderately sized programs.
-Towards this, in this work, we pursue the goal of desigining a faster algorithm, and propose an FPRAS which runs in time $\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.
+In a recent breakthrough by Arenas, Croquevielle, Jayaram and Riveros (Arenas et al., 2019, 2021), #NFA was finally shown to admit fully polynomial time randomized approximation scheme (FPRAS).
+One would assume that the discovery of FPRAS for #NFA would lead to design of scalable tools for all the different applications outlined above.
+Such has, however, not been the case and the primary hurdle to achieving practical adoption has been the prohibitively large time complexity of the proposed FPRAS– it takes time $\widetilde{O}\left(m^{17}n^{17}\cdot\frac{1}{\epsilon^{14}}\cdot\log(1/\delta)\right)$, where $\widetilde{O}$ hides the poly-logarithmic multiplicative factors in $m$ and $n$.
+It is worth highlighting that for most applications, the reduction to #NFA instance often only incurs a blow up that is a low-degree polynomial (linear or quadratic) factor larger than the size of the original problem.
+This means that the bottleneck to solving the original counting problem is really the high runtime complexity involved in the (approximate) counting algorithm and not the reduction.
+In this work, we pursue the goal of designing a faster algorithm, and propose an FPRAS which runs in time $\widetilde{O}((m^{2}n^{10}+m^{3}n^{6})\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.
 The reliance of both FPRASes on the inter-reducibility of sampling and counting necessitates a more informative measure, such as the number of samples maintained for each state.
-The scheme proposed in  (Arenas et al., 2021) maintains $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ samples per state while the scheme proposed in this work maintains only $O(\frac{n^{4}}{\epsilon^{2}})$ samples, which is independent of $m$ and with a significant reduction in it’s dependence on $n$ and $\epsilon$.
+The scheme proposed in  (Arenas et al., 2021) maintains $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ samples per state while the scheme proposed in this work maintains only $O(\frac{n^{4}}{\epsilon^{2}})$ samples, which is independent of $m$ and with a significant reduction in its dependence on $n$ and $\epsilon$.
 1.1. Overview of our FPRAS
 At a high-level, our approach exploits the tight connection between approximate counting and the related problem of (almost) uniform generation.
-In the context of an NFA $\mathcal{A}$ and a length $n$, the almost uniform generation problem asks for a randomized algorithm (generator) $G$, such that, on each invocation, the probability with which $G$ returns any given string $x\in\mathcal{L}(\mathcal{A}_{n})$ lies in the range $[\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}-\epsilon,\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}+\epsilon]$; here $\epsilon\in(0,1)$ is a parameter.
-Such a random generator is a polynomial time generator if its running time is polynomial in $|\mathcal{A}|$ and $n$.
+In the context of an NFA $\mathcal{A}$ and a length $n$, the almost uniform generation problem asks for a randomized algorithm (generator) $G$, such that, on each invocation, the probability with which $G$ returns any given string $x\in\mathcal{L}(\mathcal{A}_{n})$ lies in the range $[\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}(1-\epsilon),\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}(1+\epsilon)]$; here $\epsilon\in(0,1)$ is a parameter.
+Such a random generator is a polynomial time generator if its running time is polynomial in $|\mathcal{A}|$, $n$ and $1/\epsilon$.
 The existence of a polynomial time almost uniform generator for self-reducible (Jerrum et al., 1986) problems implies the existence of an FPRAS, and vice versa.
-Input : NFA $\mathcal{A}$ with $m$ states $Q$, $n\in\mathbb{N}$ (in unary)
+This tight connection between counting and sampling was also the high-level idea behind the work of (Arenas et al., 2021).
+Consequently, our algorithm fits the template procedure of Fig.
+1, also outlined in (Arenas et al., 2021).
+For technical convenience, the algorithm unrolls the automaton $\mathcal{A}$ into an acyclic graph [^1] The assumption in the template algorithm of Fig.
+1 that $\mathcal{A}$ has only 1 accepting state is without loss of generality.
+with $n$ levels; each level $\ell$ containing the $\ell^{\text{th}}$ copy ${q}^{\ell}$ of each state $q$ of $\mathcal{A}$.
+A key component in the algorithm is to incrementally compute an estimate $\mathsf{N}({q}^{\ell})$ of the size of the set $\mathcal{L}({q}^{\ell})$ of words of length $\ell$ that have a run ending in state $q$.
+Next, in order to inductively estimate $\mathsf{N}({q}^{\ell})$, we in turn rely on the estimates of sets $\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}^{\ell-1})$ and $\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}^{\ell-1})$; here $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is a ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor state of $q$ (${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\Sigma=\{0,1\}$), i.e., $(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta$ is a transition of $\mathcal{A}$.
+These estimates can be put to use effectively because:
+$\mathcal{L}({q}^{\ell})=\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}\}\uplus\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}.$
+Figure 1. Algorithm Template: FPRAS for #NFA
+Input : NFA $\mathcal{A}$ with $m$ states $Q$ and a single final state $q_{F}\in Q$, $n\in\mathbb{N}$ (in unary)
 1 Unroll the automaton $\mathcal{A}$ into an acyclic graph $\mathcal{A}_{\textsf{unroll}}$ by making $n+1$ copies $\{q^{\ell}\}_{\ell\in[0,n],q\in Q}$ of the states $Q$ and adding transitions between immediate layers foreach $\ell\in\{0,\ldots,n\},q\in Q$ do
-2    Compute the estimate $\mathsf{N}({q}^{\ell})$ using prior estimates and samples $\{\mathsf{N}({q}^{i}),\mathsf{S}({q}^{i})\}_{i<\ell,p\in Q}$ Compute a uniform set of samples $\mathsf{S}({q}^{\ell})$ using $\mathsf{N}({q}^{\ell})$ and the prior estimates and samples $\{\mathsf{N}({q}^{i}),\mathsf{S}({q}^{i})\}_{i<\ell,p\in Q}$
-return $\mathsf{N}({F}^{\ell})$
-Algorithm 1 Algorithm Template: FPRAS for #NFA
-This tight connection between counting and sampling was also the high-level idea behind the work of (Arenas et al., 2021).
-Consequently, our algorithm fits the template procedure of Algorithm 1, also outlined in (Arenas et al., 2021).
-For technical convenience, the algorithm unrolls the automaton $\mathcal{A}$ into an acyclic graph with $n$ levels; each level $\ell$ containing the $\ell^{\text{th}}$ copy ${q}^{\ell}$ of each state $q$ of $\mathcal{A}$.
-A key component in the algorithm is to incrementally compute an estimate $\mathsf{N}({q}^{\ell})$ of the size of the set $\mathcal{L}({q}^{\ell})$ of words of length $\ell$ that have a run ending in state $q$.
-Next, in order to inductively estimate $\mathsf{N}({q}^{\ell})$, we in turn rely on the estimates of sets $\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ (${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\Sigma=\{0,1\}$); here $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is a ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor state of $q$, i.e., $(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta$ is a transition of $\mathcal{A}$.
-These estimates can be put to use effectively because:
-$\mathcal{L}({q}^{\ell})=\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}\}\cup\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}.$
-We remark that $\mathcal{L}({p}^{\ell-1})$ may not be disjoint for different predecessor states $p$.
+2    Compute the estimate $\mathsf{N}({q}^{\ell})$ using prior estimates and samples $\{\mathsf{N}({p}^{i}),\mathsf{S}({p}^{i})\}_{i<\ell,p\in Q}$ Compute a uniform set of samples $\mathsf{S}({q}^{\ell})$ using $\mathsf{N}({q}^{\ell})$ and the prior estimates and samples $\{\mathsf{N}({p}^{i}),\mathsf{S}({p}^{i})\}_{i<\ell,p\in Q}$
+return $\mathsf{N}({q_{F}}^{n})$
+We remark that the sets $\{\mathcal{L}({p}^{\ell-1})\}_{p}$ may not be disjoint for different predecessor states $p$.
 Hence, we cannot estimate $\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ by merely summing the individual estimates $\{\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})\,|\,(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta\}$.
@@
 We then sample the final character in proportion to the size of these estimates.
-In a similar recursive manner, we can sample strings, character by character.
+Similarly, we can sample strings, character by character.
 That is, we can sample strings by progressively extending suffixes of $w$ backwards!
 Of course, we must also account for the errors that propagate in the estimates of $\mathsf{N}({q}^{\ell})$.
-Our algorithm also relies on the self-reducible union property, and therefore, the basic structure of our algorithm is similar to that of (Arenas et al., 2021) on high-level.
+Since our algorithm also relies on the self-reducible union property, the basic structure of our algorithm is similar to that of (Arenas et al., 2021) on high-level.
 There are, however, significant and crucial technical differences that allow us to obtain an FPRAS with significantly lower time complexity.
-We highlight these differences below.
-As a first step, we highlight the two key invariants underpinning the algorithm and ensuing analysis in (Arenas et al., 2021):
-  1. (1)
-    The following condition (Inv-ACJR) holds for every level $\ell$ (here, $\kappa=\frac{nm}{\epsilon}$), with high probability.
-    (Inv-ACJR) $\begin{array}[]{l}\xi(\ell)\triangleq\forall q{\in}Q,\forall P{\subseteq}Q,\big{|}\frac{|\mathsf{S}({q}^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathsf{S}({q}^{\ell})|}-\frac{|\mathcal{L}(q^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}\big{|}\leq\frac{1}{\kappa^{3}}\end{array}$
-  2. (2)
-    $\mathsf{S}({q}^{\ell})$ is close in $(L_{\infty}$ distance)111 In (Arenas et al., 2021), the invariant is stated in a different, but equivalent, formulation: conditioned on $\xi(\ell)$, the set $\mathsf{S}({q}^{\ell})$ follows the distribution of a multi-set obtained by independently sampling every element of $\mathcal{L}(q^{\ell})$ uniformly at random, with replacement.
+We highlight these differences below by first highlighting the two key properties underpinning the algorithm and ensuing analysis in (Arenas et al., 2021):
+  1. (ACJR-1).
+    For every level $\ell$, the following condition holds with high probability (here, $\kappa=\frac{nm}{\epsilon}$):
+    $\begin{array}[]{l}\xi(\ell)\triangleq\forall q{\in}Q,\forall P{\subseteq}Q,\ \left|\frac{|\mathsf{S}({q}^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathsf{S}({q}^{\ell})|}-\frac{|\mathcal{L}(q^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}\right|\leq\frac{1}{\kappa^{3}}\end{array}$
+  2. (ACJR-2).
+    $\mathsf{S}({q}^{\ell})$ is close in $(L_{\infty}$ distance)222 In (Arenas et al., 2021), the invariant is stated in a different, but equivalent, formulation: conditioned on $\xi(\ell)$, the set $\mathsf{S}({q}^{\ell})$ follows the distribution of a multi-set obtained by independently sampling every element of $\mathcal{L}(q^{\ell})$ uniformly at random, with replacement.
 to the distribution of multi-set obtained by independently sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
 In contrast, the FPRAS we propose maintains the following two weaker invariants:
-  1. (1)
+  1. (Inv-1).
     For every state $q$ and level $\ell$, the event that $\mathsf{N}({q}^{\ell})\in(1\pm\frac{\epsilon}{2n^{2}})|\mathcal{L}(q^{\ell})|$, which we denote as $\texttt{AccurateN}_{q,\ell}$, happens with high probability.
-  2. (2)
-    The distribution of $\mathsf{S}({q}^{\ell})$ is close, in total variational distance, to the distribution of a multi-set constructed by sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
+  2. (Inv-2).
+    The distribution of $\mathsf{S}({q}^{\ell})$ is close, in total variation distance, to the distribution of a multi-set constructed by sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
 Two remarks are in order.
@@
 In contrast, for $\texttt{AccurateN}_{q,\ell}$ to hold with a probability of $1-\eta$, our FPRAS only maintains $O\left(\frac{n^{4}}{\epsilon^{2}}\right)\log\eta^{-1}$ samples in $\mathsf{S}({q}^{\ell})$.
-Further, a union bound over only $O(mn)$ events is required, thus setting $\eta^{-1}=O(mn)$ is adequate.
-The time complexity of an algorithm that follows the template outlined in Algorithm 1 (including both the FPRAS of (Arenas et al., 2021) as well as the FPRAS developed in this paper), can be expressed as $O(n\cdot(m\cdot n)\cdot(m^{2}n)\cdot|\mathsf{S}({q}^{\ell})|^{2})$ where the first $n$ from the length of every string, the factor $(m\cdot n)$ corresponds to the upper bound on the number of states in the unrolled automaton, and the factor $(m^{2}n)$ corresponds to the time taken for membership check.
+Furthermore, a union bound over only $O(mn)$ events is required, thus setting $\eta^{-1}=O(mn)$ is adequate.
+The time complexity of an algorithm that follows the template outlined in Fig.
+1 (including both the FPRAS of (Arenas et al., 2021) as well as the FPRAS developed in this paper), has a quadratic depends on $|\mathsf{S}({q}^{\ell})|$.
 Accordingly, it is instructive to compare the bound of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ for $|S(q^{\ell})|$ in  (Arenas et al., 2021) to $O(\frac{n^{4}}{\epsilon^{2}})$ in our approach.
 Remarkably, the number of samples for every state in our approach is independent of $m$, the size of the automaton.
+Organization.
+In Section 2 we recall some helpful background as well as introduce a key concept (distribution entanglement) useful for the rest of the paper.
+In Section 3, we describe our algorithm, together with two auxiliary algorithms and state their formal correctness guarantees, whose analysis we present in Section 4.
+We conclude in Section 5.
 2. Preliminaries
+NFA and words.
 We consider the binary alphabet $\Sigma=\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$.
@@
 For simplicity, this paper assumes a singleton set of accepting states, i.e., $F=\{q_{F}\}$ for some $q_{F}\in Q$.
-We introduce additional notation for our algorithms.
-For $q\in Q$ and $b\in\Sigma$, the $b$-predecessors of $q$ are:
-$\textsf{Pred}(q,b)=\{p\,|\,(p,b,q)\in\Delta\}.$
-For each $q\in Q$ and $1\leq\ell\leq n$, $\mathcal{L}(q^{\ell})$ denotes the set of strings $w$ of length $\ell$ with a run on $\mathcal{A}$ ending in $q$.
-The rest of the paper focuses on the automaton $\mathcal{A}=(Q,I,\Delta,F)$, and a number $n$ in unary, using $m$ to denote $|Q|$.
+For $q\in Q$ and ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\Sigma$, the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessors of $q$ are given by the set $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})=\{p\,|\,(p,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta\}.$ Given automaton $\mathcal{A}$ and a number $n$ in unary, we construct the unrolled automaton $\mathcal{A}_{\textsf{unroll}}$, and assume that all states in $\mathcal{A}_{\textsf{unroll}}$ are reachable from the initial state.
+We use the notation ${q}^{\ell}$ to denote the $\ell^{\text{th}}$ copy of state $q\in Q$.
+For each state $q$ and level $0\leq\ell\leq n$, we use $\mathcal{L}({q}^{\ell})$ to denote the the set of all distinct strings $w$ of length $\ell$ for which there is a path (labeled by $w$) from the starting state to $q$.
+Total Variation Distance.The total variation distance between two random variables $X,Y$ over domain $\Omega$ is defined as:
+$\mathsf{TV}(X,Y)=\sum_{\omega\in\Omega}\Pr[X=\omega]-\min(\Pr[X=\omega],\Pr[Y=\omega]).$
+2.1. A New Notation: Distribution Entanglement
+We introduce a new notation, referred to as distribution entanglement, to argue about the probability of events for cases when the distribution of a certain random variable were to mimic distribution of another random variable.
+For an event $\mathcal{E}$ and random variables $X$ and $Y$, we use ${\mathsf{Pr}}_{X}\!\left[\mathcal{E};{Y}\right]$ to denote the probability of $\mathcal{E}$ when $X$ has distribution entangled to $Y$, which is defined as follows:
+${\mathsf{Pr}}_{X}\!\left[\mathcal{E};{Y}\right]=\sum_{\omega\in\Omega}{\mathsf{Pr}}\!\left[Y=\omega\right]\cdot{\mathsf{Pr}}\!\left[\mathcal{E}|X=\omega\right].$
+where $\Omega$ is the support of $Y$.
+The above notion is well-defined only when it is the case for all $\omega\in\Omega$, we have $\Pr[X=\omega]>0$.
+Likewise, we can also extend this notion for conditioned events.
+That is, for events $\mathcal{E}$, $\mathcal{F}$ and random variables $X$ and $Y$, the conditional probability of $\mathcal{E}$ given $\mathcal{F}$ when $X$ has distribution entangled to $Y$is defined as follows:
+${\mathsf{Pr}}_{X}\!\left[\mathcal{E}|\mathcal{F};{Y}\right]=\sum_{\omega\in\Omega}{\mathsf{Pr}}\!\left[Y=\omega\right]\cdot{\mathsf{Pr}}\!\left[\mathcal{E}|\mathcal{F}\cap X=\omega\right].$
+Whenever $X$ is clear from the context, we will omit it from the notation.
 3. A Faster FPRAS
-Given automaton $\mathcal{A}$ and a number $n$ in unary, we construct the unrolled automaton $\mathcal{A}_{\textsf{unroll}}$.
-We use the notation ${q}^{\ell}$ to denote the $\ell^{\text{th}}$ copy of state $q\in Q$.
-We use $\mathcal{L}({q}^{\ell})$ to denote the the set of all distinct strings $w$ of length $\ell$ for which there is a path (labeled by $w$) from the starting state to $q$ (i.e., those strings that go to state ${q}^{\ell}$ in $\mathcal{A}_{\textsf{unroll}}$).
-We will use $\mathsf{N}({q}^{\ell})$ to denote an estimate of $|\mathcal{L}({q}^{\ell})|$.
-In Section 3.1, we formally present our FPRAS for #NFA, together with the statement of its correctness and running time (Theorem 3).
-Our FPRAS Algorithm 4 calls two subroutines Algorithm 3 and Algorithm 2, each of which are presented next, together with their informal descriptions.
-In later sections, we present the formal proofs of their correctness.
-3.1. Description of the Algorithms
-Algorithm 4 presents the main procedure for approximating the size of $\mathcal{L}(\mathcal{A})_{n}$, where $\mathcal{A}$ is the NFA and $n$ is the parameter that is given as input.
-The algorithm works in a dynamic programming fashion and maintains, for each state $q$ and each length $1\leq\ell\leq n$, a sampled subset $\mathsf{S}({q}^{\ell})$ of $\mathcal{L}({q}^{\ell}))$.
-As a key step, it uses the sampled sets corresponding to $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}})$ and $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}})$ to construct the desired sampled subset of $\mathcal{L}({q}^{\ell}))$.
-We first describe the two subroutines the $\texttt{AppUnion}{()}$ and $\mathsf{sample}()$.
-We start with describing the subroutines first.
+In this section we formally present our FPRAS for #NFA, together with the statement of its correctness and running time (Theorem 3).
+Our FPRAS Algorithm 3 calls two subroutines AppUnion (Algorithm 1) and sample (Algorithm 2), each of which are presented next, together with their informal descriptions and formal statements of correctness.
+Algorithm 3 presents the main procedure for approximating the size of $\mathcal{L}(\mathcal{A}_{n})$, where $\mathcal{A}$ is the NFA and $n$ is the parameter that is given as input.
+The algorithm works in a dynamic programming fashion and maintains, for each state $q$ and each length $1\leq\ell\leq n$, a sampled subset $\mathsf{S}({q}^{\ell})$ of $\mathcal{L}({q}^{\ell})$.
+As a key step, it uses the sampled sets corresponding to $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}})$ and $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}})$ to construct the desired sampled subset of $\mathcal{L}({q}^{\ell})$.
+We first describe the two subroutines the AppUnion and sample.
+3.1. Approximating Union of Sets
+AppUnion (Algorithm 1) takes parameters $\epsilon,\delta$ and inputs $\epsilon_{\textsf{sz}}$ and (access to) sets $T_{1},\dots,T_{k}$.
+For each set $T_{i}$ the access is given in the form of a membership oracle $O_{i}$, a subset $S_{i}$ (obtained by sampling $T_{i}$ with replacements) presented as a list, and an estimate $\textsf{sz}_{i}$ of the size of $T_{i}$.
+Using these, Algorithm 1 outputs an $(\epsilon,\delta)$ estimate of the size of the union $\cup_{i=1}^{k}T_{i}$.
+The precise algorithm presented here represents a modification of the classic Monte Carlo-based scheme due to Karp and Luby (Karp and Luby, 1985).
+The proof of its correctness also shares similarities with (Karp and Luby, 1985) and we summarize the intuition here.
+For a $\sigma\in\cup_{i=1}^{k}T_{i}$, let $L(\sigma)$ denote the smallest index $i\in\{1,2,\ldots,k\}$ such that $\sigma\in T_{i}$.
+Now, let $\mathcal{U}_{\textsf{unique}}$ be the set of all pairs $(\sigma,L(\sigma))$ and let $\mathcal{U}_{\textsf{multiple}}$ be the set of all $(\sigma,i)$ pairs such that $\sigma\in T_{i}$.
+Consequently, the cardinality of the set $\mathcal{U}_{\textsf{unique}}$ is $|\cup_{i=1}^{k}T_{i}|$, while the cardinality of $\mathcal{U}_{\textsf{multiple}}$ is $\sum_{i=1}^{k}|T_{i}|$.
+By drawing sufficiently-many ($t$, to be precise) samples from $\mathcal{U}_{\textsf{multiple}}$ and assessing the fraction belonging to $\mathcal{U}_{\textsf{unique}}$, one can estimate $|\mathcal{U}_{\textsf{unique}}|/|\mathcal{U}_{\textsf{multiple}}|$, which when multiplied by $|\mathcal{U}_{\textsf{multiple}}|$ provides an estimate for the size of $\mathcal{U}_{\textsf{unique}}$, which is the desired output.
 1 function AppUnion ${}_{(\epsilon,\delta)}$($\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$)
-2     $m\leftarrow\lceil\frac{\sum_{i}\textsf{sz}_{i}}{\max_{i}\textsf{sz}_{i}}\rceil$ $t\leftarrow\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot m}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ $Y\leftarrow 0$ repeat $t$ times:
-3          With prob. $p_{i}=\frac{\textsf{sz}_{i}}{\sum_{i}\textsf{sz}_{i}}$, sample $i\in\{1,\ldots,k\}$ if $S_{i}\neq\varnothing$ then
-4               $\sigma\leftarrow S_{\ell}\!\boldsymbol{\cdot}\!\mathsf{dequeue}()$
-5         else
-6               break
-7         if $\not\exists j<i,$ such that $O_{j}\!\boldsymbol{\cdot}\!\mathsf{contains}(\sigma)$ then  $Y\leftarrow Y+1$
-8    return $\frac{Y}{t}\cdot{\sum_{i}\textsf{sz}_{i}}$
+2     $m\leftarrow\left\lceil\frac{\sum_{j}\textsf{sz}_{j}}{\max_{j}\textsf{sz}_{j}}\right\rceil$ $t\leftarrow\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot m}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ $Y\leftarrow 0$ repeat $t$ times:
+3          Sample $i\in\{1,\ldots,k\}$ with probability $p_{i}=\frac{\textsf{sz}_{i}}{\sum_{j}\textsf{sz}_{j}}$ if $S_{i}\neq\varnothing$ then  $\sigma\leftarrow S_{i}\!\boldsymbol{\cdot}\!\mathsf{deque}(){}$ else  break if $\neg(\exists j<i,O_{j}\!\boldsymbol{\cdot}\!\mathsf{contains}(\sigma))$ then  $Y\leftarrow Y+1$
+4    return $\frac{Y}{t}\cdot({\sum_{i}\textsf{sz}_{i}})$
 Algorithm 1 Approximating Union of Sets
+After sampling an index $i$, in Line 1 an element is chosen (and removed) from the sampled list $S_{i}$.
+In the case when the set $S_{i}$ is constructed by uniformly selecting elements from $T_{i}$, this step mimics drawing a random sample from the actual set $T_{i}$.
+Initially, each set (more precisely, list) $S_{i}$ is guaranteed to have a size of at least $\texttt{thresh}=24\cdot\frac{(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\cdot\log(\frac{4k}{\delta})$, surpassing the expected number of samples required from $T_{i}$ during the $t$ iterations of the loop.
+If the algorithm ever needs more samples than $|S_{i}|$ during the $t$ iterations, it counts as an error, but the probability of this occurrence will be demonstrated to be very low.
+Finally, Line 1 verifies whether the sample from $\mathcal{U}_{\textsf{multiple}}$ belongs to $\mathcal{U}_{\textsf{unique}}$ by asking a membership question to the oracle $O_{i}$.
+If all the sampled sets $S_{i}$’s are constructed uniformly, then the variable $Y$ tallies the number of samples from $\mathcal{U}_{\textsf{unique}}$, and after $t$ iterations, $Y/t$ provides a estimation for $|\mathcal{U}_{\textsf{unique}}|/|\mathcal{U}_{\textsf{multiple}}|$.
+The final output is the product of this value with $(\sum_{i}\textsf{sz}_{i})$ (which is the estimate for $|\mathcal{U}_{\textsf{multiple}}|$).
+We next present (in Theorem 1) the formal correctness statement of this algorithm using distribution entanglement (see Section 2.1).
+For this, we set the source random variable $X$ to be the one that corresponds to the product distribution of the input sampled sets $S_{1},\ldots,S_{k}$, while the target random variable $Y$ corresponds to the product distribution of ${\mathsf{U}}_{1},\ldots,{\mathsf{U}}_{k}$, where, ${\mathsf{U}}_{i}$, is the random variable that obeys the distribution obtained when constructing a subset of $T_{i}$ (of at least thresh many elements) by uniformly picking each element of $T_{i}$.
+Since the random variables $S_{1},\ldots,S_{k}$ are clear from context, we will use ${\mathsf{Pr}}\!\left[\mathcal{E};{{\mathsf{U}}_{1,\ldots,k}}\right]$ to denote the probability of event $\mathcal{E}$ in the resulting entanglement distribution.
+For well definedness, we require that ${\mathsf{Pr}}\!\left[S_{i}=\omega\right]>0$ for every subset $\omega\subseteq T_{i}$ (with $\omega\geq\texttt{thresh}$), and will ensure this each time we invoke this statement.
+Theorem 1 ().
+Let $\epsilon,\delta>0$, and let $\Omega$ be some set.
+Let $T_{1},T_{2},\ldots,T_{k}\subseteq\Omega$ be sets with membership oracles $O_{1},\ldots,O_{k}$ respectively.
+Let $\epsilon_{\textsf{sz}}\geq 0$ and let $\textsf{sz}_{1},\ldots,\textsf{sz}_{k}\in\mathbb{N}$ be such that for every $i\leq k$, we have $\frac{\left|T_{i}\right|}{(1+\epsilon_{\textsf{sz}})}\leq\textsf{sz}_{i}\leq(1+\epsilon_{\textsf{sz}})\left|T_{i}\right|$.
+Let $\texttt{thresh}=24\cdot\frac{(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\cdot\log(\frac{4k}{\delta})$.
+For each $i\leq k$, let $S_{i}$ be some sequence of samples (with possible repetitions), obtained according to some distribution such that $|S_{i}|\geq\texttt{thresh}$ and for every $\omega\subseteq T_{i}$ with $|\omega|\geq\texttt{thresh}$, ${\mathsf{Pr}}\!\left[S_{i}=\omega\right]>0$.
+Then, on input $\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$, the algorithm AppUnion with parameters $\epsilon,\delta$ (Algorithm 1) outputs sz that satisfies:
+${\mathsf{Pr}}\!\left[\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}\leq\left|\textsf{sz}\right|<(1+\epsilon)(1+\epsilon_{\textsf{sz}})|\cup_{i=1}^{k}T_{i}|\,;{{\mathsf{U}}_{1,\ldots,k}}\right]>1-\delta.$
+The algorithm makes $O\left(k\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot\frac{1}{\epsilon^{2}}\cdot\log\big{(}\frac{k}{\delta}\big{)}\right)$ number of membership calls to the oracles $O_{1},\ldots,O_{k}$ in the worst case and the rest of the computation takes a similar amount of time.
+3.2. Sampling Subroutine
+[…]
 1 function $\ell$, ${P}^{\ell}$, $w$, $\varphi$,$\beta$, $\eta$)
@@
 5          foreach ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$ do
-6               ${P}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\bigcup_{p\in{P}^{\ell}}\textsf{Pred}(p,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ // From the above set, remove all states that are not reachable from the initial state.
 […]
-8
 7         $\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}\leftarrow\frac{\displaystyle\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}{\displaystyle\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}$ w.p.
@@
 Algorithm 2 Sampling subroutine for $\mathcal{A}_{\textsf{unroll}}$, length $n$
-The AppUnion subroutine.
-AppUnion ( Algorithm 2) takes as input $\epsilon,\delta$ and $\epsilon_{\textsf{sz}}$ and assumes access to sets $T_{1},\dots,T_{k}$.
-For each set $T_{i}$ the access is given in the form of membership oracle $O_{i}$, a subset $S_{i}$ (obtained by sampling $T_{i}$ with replacements), and an estimate $\textsf{sz}_{i}$ of the size of $T_{i}$.
-Using these, Algorithm 2 outputs an $(\epsilon,\delta)$ estimate of the size of the union $\cup_{i=1}^{k}T_{i}$.
-The algorithm presented herein represents a modification of classical Monte Carlo-based scheme due to Karp and Luby (Karp and Luby, 1985).
-The proof methodology also shares similarities with  (Karp and Luby, 1985).
-For an arbitrary $\sigma\in\cup_{i=1}^{k}T_{i}$, let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$.
-And let set $\mathcal{U}_{1}$ be the set of all the $(\sigma,L(\sigma))$ pairs and let $\mathcal{U}_{2}$ be the set of all $(\sigma,i)$ pairs such that $\sigma\in T_{i}$.
-Consequently, the cardinality of the set $\mathcal{U}_{1}$ is $|\cup_{i=1}^{k}T_{i}|$, and that of $\mathcal{U}_{2}$ is $\sum_{i=1}^{k}|T_{i}|$.
-By drawing a sufficient number of samples from $\mathcal{U}_{2}$ and assessing the fraction belonging to $\mathcal{U}_{1}$, one can estimate $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$, which when multiplied by $|\mathcal{U}_{2}|$ provides a reliable estimate for the size of $\mathcal{U}_{1}$.
-In the algorithm we take $t$ samples from $\mathcal{U}_{2}$.
-This involves initially sampling a set $T_{i}$ proportional to the size of the set (in line 2 ) and subsequently uniformly drawing a sample from $T_{i}$ (in line 2).
-As the exact sizes of the sets $T_{i}$ are unknown , estimates $\textsf{sz}_{i}$ are utilized for proportional sampling (in line 2), introducing an additional factor of error in the final estimate of $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$.This error can be mitigated by adjusting the value of $t$ accordingly.
-Upon sampling an $i$ in Line 2, an element is chosen from the sample of the sets $S_{i}$.
-This mimics a random sample from $T_{i}$, given that $S_{i}$ already contains uniform samples from $T_{i}$.
-For each sampled index $i$ , a new element from $S_{i}$ is used and removed (Line 2).
-Initially, $S_{i}$ are sets are guaranteed to have a size of at least thresh, surpassing the expected number of samples required from $T_{i}$ during the $t$ iterations of the loop.
-If the algorithm ever needs more samples than $|\mathcal{U}_{i}|$ during the $t$ iterations, it counts as an error, but the probability of this occurrence is demonstrated to be very low.
-Finally, line 2 verifies whether the sample from $\mathcal{U}_{2}$ belongs to $\mathcal{U}_{1}$.
-It uses the oracles to the sets to check this.
-The variable $Y$ tallies the number of samples from $\mathcal{U}_{1}$, and after $t$ iterations, $Y/t$ provides a estimation for $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$.
-The final output is the product of this value with $(\sum_{i}\textsf{sz}_{i})$, the estimate for $|\mathcal{U}_{2}|$.
-Theorem 1 characterizes the formal guarantee of Algorithm 2.
-Theorem 1.
-Let $\epsilon,\delta>0$, and let $\Omega$ be some set.
-Let $T_{1},T_{2},\ldots,T_{k}\subseteq\Omega$ be sets with membership oracles $O_{1},\ldots,O_{k}$ respectively.
-Let $\epsilon_{\textsf{sz}}\geq 0$ and let $\textsf{sz}_{1},\ldots,\textsf{sz}_{k}\in\mathbb{N}$ be such that for every $i\leq k$, we have $\frac{\left|T_{i}\right|}{(1+\epsilon_{\textsf{sz}})}\leq\textsf{sz}_{i}\leq(1+\epsilon_{\textsf{sz}})\left|T_{i}\right|$.
-For each $i\leq k$, let $S_{i}$ be a sequence of samples (with possible repetitions) obtained by repeatedly sampling elements from $T_{i}$ uniformly with replacement such that $|S_{i}|\geq\texttt{thresh}$, where $\texttt{thresh}=24\cdot\frac{(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\cdot\log(\frac{4k}{\delta})$ Then, on input $\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$, the algorithm AppUnion with parameters $\epsilon,\delta$ (Algorithm 2) outputs sz that satisfies:
-${\mathsf{Pr}}\!\left[\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}\leq\left|\textsf{sz}\right|<(1+\epsilon)(1+\epsilon_{\textsf{sz}})|\cup_{i=1}^{k}T_{i}|\right]>1-\delta.$
-The worst case number of calls to the oracles $O_{i}$ made by the algorithm is $O\left(k\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot\frac{1}{\epsilon^{2}}\cdot\log\big{(}\frac{k}{\delta}\big{)}\right)$ and the rest of the computation takes a similar amount of time.
-[…]
 We next tread towards formally characterizing the guarantee of Algorithm 2. To this end, we will introduce few notations:
-  1. (1)
   • For each $1\leq i\leq\ell$ and each $q\in Q$, let $\texttt{AccurateN}_{i,q}$ be the event that $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|$.
 Also, let $\texttt{AccurateN}_{i}=\bigcap\limits_{q\in Q}\texttt{AccurateN}_{i,q}$ and $\texttt{AccurateN}_{\leq i}=\bigcap\limits_{1\leq j\leq i}\texttt{AccurateN}_{j}$.
-  2. (2)
-    For a state ${q}^{\ell}$, we will use $\mathsf{U}({q}^{\ell})$ to denote the random variable denoting the sequence of samples (with possible repetitions) constructed by repeatedly sampling $|\mathsf{S}({q}^{\ell})|$ number of elements uniformly from $\mathcal{L}({q}^{\ell})$.
-Our algorithms compute sequences $\mathsf{S}({q}^{\ell})$ of elements (strings) sampled from $\mathcal{L}({q}^{\ell})$.
-For arguing about correctness of our algorithms, we will in turn argue that the distribution of $\mathsf{S}({q}^{\ell})$ would be “similar" to that of $\mathsf{U}({q}^{\ell})$.
-Towards this, we will often invoke the the probability space obtained by replacing the sets $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$.
-Formally, for any event $\mathcal{E},\mathcal{F}$ (in the context of Algorithm 3 and Algorithm 4) we will denote by ${\mathsf{Pr}}\!\left[\mathcal{E};{{\mathsf{U}}_{\leq i}}\right]$ (resp.
-${\mathsf{Pr}}\!\left[\mathcal{E}\;|\;\mathcal{F};{{\mathsf{U}}_{\leq i}}\right]$) the probability of the event $\mathcal{E}$ (resp.
-probability of the event $\mathcal{E}$, given event $\mathcal{F}$) when the $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 is replaced by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$.
-  3. (3)
+  • For a state ${q}^{\ell}$, we use $\mathsf{U}({q}^{\ell})$ to represent the random variable denoting the sequence of samples (with possible repetitions) constructed by repeatedly sampling $|\mathsf{S}({q}^{\ell})|$ number of elements uniformly from $\mathcal{L}({q}^{\ell})$.
+Let ${\sf S}_{\leq i}$ be the random variable corresponding to the product of $\{\mathsf{S}({q}^{j})\}_{q\in Q,0\leq j\leq i}$.
+Also, let ${\mathsf{U}}_{\leq i}$ represent the random variable obtained by taking the ordered profuct of all $\{\mathsf{U}({q}^{j})\}_{q\in Q,0\leq j\leq i}$.
+In our algorithms, we will use ${\mathsf{Pr}}\!\left[\mathcal{E};{{\mathsf{U}}_{\leq i}}\right]$ (resp.
+${\mathsf{Pr}}\!\left[\mathcal{E}|\mathcal{F};{{\mathsf{U}}_{\leq i}}\right]$) to denote the probability (resp.
+conditional probability) of $\mathcal{E}$ (resp.
+$\mathcal{E}$ conditioned on $\mathcal{F}$) when ${\sf S}_{\leq i}$ is distribution entangled to ${\mathsf{U}}_{\leq i}$.
   • Consider the random trial corresponding to the call to the procedure $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$ for $\ell>0$ and let $\mathtt{ret}_{\textnormal{{sample}}}$ denote the random variable representing the return value thus obtained.
@@
 Let us now define some events of interest.
-  4. (4)
-    For each $1\leq i\leq\ell$ and ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$, let us denote by $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ the event that the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}^{\text{th}}$ call to AppUnion corresponding to level $i$ returns a value $\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ that satisfies 1($\frac{1}{(1+\beta)^{i}}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|\leq\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq(1+\beta)^{i}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|$.
+  • For each $1\leq i\leq\ell$ and ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$, by $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ we denote the event that the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}^{\text{th}}$ call to AppUnion corresponding to level $i$ returns a value $\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ that satisfies 1($\frac{1}{(1+\beta)^{i}}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|\leq\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq(1+\beta)^{i}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|$.
 Here, $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$.
 For ease of notation, we also define $\texttt{Accurate}\texttt{AppUnion}=\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$
+  • Let $\mathtt{Fail}_{1}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned because $\varphi>1$ at the time of return (Line 2).
+Let $\mathtt{Fail}_{2}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned at Line 2.
+Finally, let $\mathtt{Fail}=\mathtt{Fail}_{1}\uplus\mathtt{Fail}_{2}$.
+The following presents the correctness of Algorithm 2, while its proof is presented in Section 4.2.
+Theorem 2 ().
+Let $\gamma_{0}=\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})},\beta\leq\frac{1}{4\cdot n^{2}}$ and $\{\mathsf{S}({p}^{i})\}_{1\leq i<\ell,p\in Q}$ satisfy
+ $\forall\;1\leq i\leq\ell,\forall p\in Q,\;\;|\mathsf{S}({p}^{i})|\geq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right).$
+Then for the trial corresponding to the invocation $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$ the following hold
   1. (1)
-    Let $\mathtt{Fail}_{1}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned because $\varphi>1$ at the time of return (Line 3).
-Let $\mathtt{Fail}_{2}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned at Line 3.
-Finally, let $\mathtt{Fail}=\mathtt{Fail}_{1}\uplus\mathtt{Fail}_{2}$.
-The following presents the correctness of Algorithm 3, while its proof is presented in Section 5.
-Theorem 2.
-Let $\gamma_{0},\beta$ and $\{\mathsf{S}({p}^{i})\}_{1\leq i<\ell,p\in Q}$ satisfy the following:
-$\begin{array}[]{c}\gamma_{0}=\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})},\quad\quad\beta\leq\frac{1}{4\cdot n^{2}},\quad\quad\text{ and,}\\ \forall\;1\leq i\leq\ell,\forall p\in Q,\;\;|\mathsf{S}({p}^{i})|\geq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\end{array}$
-Then for the trial corresponding to the invocation $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$, the following statements hold true
-  1. (1)
-    For each $w\in\mathcal{L}({q}^{\ell})$,
-    ${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\;\;\big{|}\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=\gamma_{0}$
+    For each $w\in\mathcal{L}({q}^{\ell})$, ${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\;\;\big{|}\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=\gamma_{0}$
   2. (2)
@@
   3. (3)
-    ${\mathsf{Pr}}\!\left[\texttt{Accurate}\texttt{AppUnion}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\frac{\eta}{2}$
-Input : NFA $\mathcal{A}$ with $m$ states, $n\in\mathbb{N}$ presented in unary
+    ${\mathsf{Pr}}\!\left[\texttt{Accurate}\texttt{AppUnion}\mid\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\frac{\eta}{2}$
+3.3. Main Algorithm
+Input : NFA $\mathcal{A}$ with $m$ states, $n\in\mathbb{N}$ in unary
 Parameters : Accuracy parameter $\varepsilon$ and confidence parameter $\delta$
-1 $\beta\leftarrow\frac{\varepsilon}{4n^{2}}$ $\eta\leftarrow\frac{\delta}{2\cdot n\cdot m}$ $\mathtt{ns}\leftarrow\frac{4096\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{4096\cdot m^{2}\cdot n^{2}\cdot\log(\varepsilon^{-2})}{\delta}\right)$ $\mathtt{xns}\leftarrow\mathtt{ns}\cdot 12\cdot(1-\frac{2}{3e^{2}})^{-1}\cdot\log(8/\eta)$ Construct the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$ from $\mathcal{A}$ input $n$ for $\ell\in\{0,\ldots,n\}\text{ and }q\in Q$ do
+1 $\beta\leftarrow\frac{\varepsilon}{4n^{2}};\eta\leftarrow\frac{\delta}{2\cdot n\cdot m}$ $\mathtt{ns}\leftarrow\frac{4096\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{4096\cdot m^{2}\cdot n^{2}\cdot\log(\varepsilon^{-2})}{\delta}\right)$ $\mathtt{xns}\leftarrow\mathtt{ns}\cdot 12\cdot(1-\frac{2}{3e^{2}})^{-1}\cdot\log(8/\eta)$ Construct the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$ from $\mathcal{A}$ for $\ell\in\{0,\ldots,n\}\text{ and }q\in Q$ do
 2     if $\ell=0$ then
-3          $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow(q\in F)$ $\big{(}1,\{\lambda\}\big{)}$ : $\big{(}0,\varnothing)$
+3          if $q=I$ then
+4               $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow(1,{\lambda})$
+5         else
+6               $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow\big{(}0,\varnothing)$
+7    else
 8          foreach ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$ do
-5          $\{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1},\ldots,p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}\}\leftarrow\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ $\beta^{\prime}=(1+\beta)^{\ell-1}-1$ $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\texttt{AppUnion}_{\beta,\eta/2}\bigg{(}\beta^{\prime},$ $(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1})),\ldots,(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1})),$ $\mathcal{A}_{\textsf{unroll}}\bigg{)}$   // The last parameter $\mathcal{A}_{\textsf{unroll}}$ is a proxy for the membership oracles required by AppUnion
-6
-7    $\mathsf{N}({q}^{\ell})\leftarrow\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ $\mathsf{S}({q}^{\ell})\leftarrow\varnothing$ repeat $\mathtt{xns}$ times:
+9               let $\{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1},\ldots,p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}\}$ be the set $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ $\beta^{\prime}\leftarrow(1+\beta)^{\ell-1}-1$ $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\texttt{AppUnion}_{\beta,\frac{\eta}{2}\cdot(1-\frac{1}{2^{n+1}})}\bigg{(}\beta^{\prime},(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1})),\ldots,(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1})),\mathcal{A}_{\textsf{unroll}}\bigg{)}$
+10         w.p.$1-\frac{\eta}{2^{n}}$ :
+11               $\mathsf{N}({q}^{\ell})\leftarrow\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$
+12         w.p.$\frac{\eta}{2^{n}}$ :
+[…]
+14          $\mathsf{S}({q}^{\ell})\leftarrow\varnothing$ repeat $\mathtt{xns}$ times:
 15               if $|\mathsf{S}({q}^{\ell})|<\mathtt{ns}$ then
@@
 17                         $\mathsf{S}({q}^{\ell})\leftarrow\mathsf{S}({q}^{\ell})\boldsymbol{\!\cdot}\!\mathsf{append}(w)$
-11
+18
 19               else break
@@
 21               let $w_{{q}^{\ell}}$ be some word in $\mathcal{L}({q}^{\ell})$ repeat $\alpha_{{q}^{\ell}}$ times:  $\mathsf{S}({q}^{\ell}){\leftarrow}\mathsf{S}({q}^{\ell})\boldsymbol{\!\cdot}\!\mathsf{append}(w_{{q}^{\ell}})$
-15
 return $\mathsf{N}({q}^{n}_{F})$
-Algorithm 4 Algorithm for Estimating $|\mathcal{L}(\mathcal{A})_{n}|$
-The main algorithm.
-Algorithm 4 first constructs (in Line 4) the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$.
+Algorithm 3 Algorithm for estimating $|\mathcal{L}(\mathcal{A}_{n})|$
+Algorithm 3 first constructs (in Line 3) the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$.
 The algorithm then goes over all the states and all levels $0\leq\ell\leq n$ to construct the sampled sets and the size estimates inductively.
-In Lines 4-4, the algorithm caters for the base case ($\ell=0$).
-In the inductive case (Lines 4-4) the algorithm computes the estimate $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$, where $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the estimate of the size of $\cup_{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}\in\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}}^{\ell-1})$.
+In Lines 3-3, the algorithm caters for the base case ($\ell=0$).
+In the inductive case (Lines 3-3) the algorithm computes the estimate $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$, where $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the estimate of the size of $\cup_{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}\in\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}}^{\ell-1})$.
 The size estimates of, and a sample subsets of the sets $\{\mathcal{L}({q}^{\ell-1})\}_{q\in Q}$ are available inductively at this point.
 Further, membership in these languages can be easily checked.
-As a result, the algorithm uses the subroutine AppUnion to compute the size of the union (in Line 4).
-Once the size estimates $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ are determined, the algorithm constructs the sampled subset of $\mathcal{L}({q}^{\ell})$ by making $\mathsf{xns}$ calls to the subroutine sample (Lines 4-4).
-Theorem 3 presents the desired correctness statement of our FPRAS  Algorithm 4, and its proof is presented in Section 6.
-Theorem 3.
+As a result, the algorithm uses the subroutine AppUnion to compute the size of the union (in Line 3), with $\mathcal{A}_{\textsf{unroll}}$ as the membership oracle as before.
+Once the size estimates $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ are determined, the algorithm constructs the sampled subset of $\mathcal{L}({q}^{\ell})$ by making $\mathsf{xns}$ calls to the subroutine sample (Lines 3-3); if fewer than $\mathtt{ns}$ many strings were sampled, the algorithm simply pads one fixed string from $\mathcal{L}({q}^{\ell})$ so that $\mathsf{S}({q}^{\ell})$ eventually has size $\mathtt{ns}$ (see Lines 3-3).
+Theorem 3 presents the desired correctness statement of our FPRAS  Algorithm 3, and its proof is presented in Section 4.3.
+Theorem 3 ().
 Given an NFA $\mathcal{A}$ with $m$ states and $n\in\mathbb{N}$ (presented in unary), Algorithm 3 returns $\mathsf{Est}$ such that the following holds:
-${\mathsf{Pr}}\!\left[\frac{|L(F^{n})|}{1+\epsilon}\leq\mathsf{Est}\leq(1+\epsilon)|L(F^{n})|\right]\geq 1-\delta$
-Moreover, the algorithm has time complexity $\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$, where the tilde hides for polynomial factors of $\log(m+n)$.
-4. Correctness of Algorithm 2
-The proof is in the same lines as that of Karp and Luby (Karp and Luby, 1985); for completeness, we present a full proof.
-Proof.
+${\mathsf{Pr}}\!\left[\frac{|\mathcal{L}(\mathcal{A}_{n})|}{1+\epsilon}\leq\mathsf{Est}\leq(1+\epsilon)|\mathcal{L}(\mathcal{A}_{n})|\right]\geq 1-\delta$
+Moreover, the algorithm has time complexity $\widetilde{O}((m^{2}n^{10}+m^{3}n^{6})\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$, where the tilde hides for polynomial factors of $\log(m+n)$.
+4. Technical Analysis
+We now present the detailed technical analyses of the algorithms presented in Section 3.
+4.1. Correctness of Algorithm 1
+The proof is in the same lines as that of Karp and Luby (Karp and Luby, 1985). For the sake of completeness, we present a full proof.
+Proof (of Theorem 1).
 Let $\mathsf{Fail}$ be the event that the output of the algorithm is not within $\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$ and $|\cup_{i=1}^{k}T_{i}|{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$.
-Clearly, we need to upper bound $\Pr[\mathsf{Fail}]$ by $\delta$.
-Note that the algorithm in each of the $t$ runs of the loop (lines 2-2) tries to draw an element from some $S_{i}$.
-So if we assume that the size of the sets $S_{i}$ for all $i$ is more than $t$, then the condition is the if (in line 2 will be satisfied and in that case the else clause in line 2 to 2 is redundant.
+We will show that ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]$ is upper-bounded by $\delta$.
+Note that the algorithm in each of the $t$ runs of the loop (Lines 1-1) tries to draw an element from some $S_{i}$.
+So if we assume that the size of the sets $S_{i}$ for all $i$ is more than $t$, then the condition in the if (in Line 1) will be satisfied and in that case the else clause in Line 1 is redundant.
 We will prove the correctness of  Algorithm 1 in two parts.
-In the first part we will prove that if the size of the sets $S_{i}$ is greater than $t$ then $\Pr[\mathsf{Fail}]\leq\frac{\delta}{2}$.
-In the second part we will prove that if the sets $S_{i}$ has size thresh (much smaller than $t$) then the probability that the algorithm will ever need the else clause in line 2 to 2 is less that $\delta/2$.
-The two parts combined proves that if the sets $S_{i}$s are of size threshthen $\Pr[\mathsf{Fail}]\leq\delta$ which is what is claimed in the theorem.
+In the first part we will prove that if the size of the sets $S_{i}$ is greater than $t$ then ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq\frac{\delta}{2}$.
+In the second part we will prove that if the sets $S_{i}$ has size thresh (much smaller than $t$) then the probability that the algorithm will ever need the else clause in Line 1 is less that $\delta/2$.
+The two parts combined proves that if the sets $S_{i}$s are of size thresh, then ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq\delta$ which is what is claimed in the theorem.
 For all $\sigma\in\cup_{i=1}^{k}T_{i}$ let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$.
 In other words, $\sigma$ is in $T_{L(\sigma)}$ and for all $i<L(\sigma)$, $\sigma\not\in T_{i}$.
-Observe $\mathcal{U}_{1}:=\left\{(\sigma,L(\sigma)\ \mid\ \sigma\in\cup_{i=1}^{k}T_{i}\right\}$ has size exactly $\left|\cup_{i=1}^{k}T_{i}\right|$.
-Furthermore, $\mathcal{U}_{2}:=\left\{(\sigma,i)\ \mid\ \sigma\in T_{i}\right\}$ has size $\sum_{i=1}^{k}|T_{i}|$.
-Part 1: If $\forall i$, $|S_{i}|\geq t$ then $\Pr[\mathsf{Fail}]\leq\frac{\delta}{2}$.
-Note that in this case the condition in the if clause in line 2 is always satisfied.
-For a run of the loop (lines 2-2), consider the lines 2 to 2.
-We say the pair $(\sigma,i)$ is sampled if in one of the iteration of the loop $i$ is sampled in line 2 and then in that same iteration $\sigma$ is obtained from $S_{i}.\mathsf{deque()}$ in the line 2.
-For any $i_{0}\in[t]$ and $\sigma_{0}\in S_{i_{0}}$, we have
+Observe $\mathcal{U}_{\textsf{unique}}:=\left\{(\sigma,L(\sigma))\ \mid\ \sigma\in\cup_{i=1}^{k}T_{i}\right\}$ has size exactly $\left|\cup_{i=1}^{k}T_{i}\right|$.
+Furthermore, $\mathcal{U}_{\textsf{multiple}}:=\left\{(\sigma,i)\ \mid\ \sigma\in T_{i}\right\}$ has size $\sum_{i=1}^{k}|T_{i}|$.
+Part 1: If $\forall i$, $|S_{i}|\geq t$ then ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq\frac{\delta}{2}$.
+If $|S_{i}|\leq t$ for all $i$, then the condition in the if clause in Line 1 is always satisfied.
+For a run of the loop (Lines 1-1), consider the Lines 1-1.
+We say the pair $(\sigma,i)$ is sampled in round $j$ if in the $j^{\text{th}}$ iteration of the loop $i$ is sampled in Line 1 and then in that same iteration $\sigma$ is obtained from $S_{i}.\mathsf{dequeue()}$ in the Line 1.
+For any $i_{0}\in\{1,\ldots,k\},\sigma_{0}\in S_{i_{0}}$ and $j_{0}\in\{1,\ldots,t\}$, we have
+[…]
+Since the choice of the set to pick $i_{0}$ is independent of the distributions of any of the $S_{i}$’s, we have,
+[…]
+Next observe that the second term in Equation (1) is ${\mathsf{Pr}}\!\left[\sigma_{0}\mbox{ is sampled in round $j_{0}$ from }S_{i_{0}};{{\mathsf{U}}_{1,\ldots,k}}\right]$.
+This term, in turn, can be computed by considering the disjoint union of the events corresponding to how many times $i_{0}$ has been chosen so far: ${\mathsf{Pr}}\!\left[\sigma_{0}\mbox{ is sampled in round $j_{0}$ from }S_{i_{0}};{{\mathsf{U}}_{1,\ldots,k}}\right]$
+$=$ $\sum\limits_{m=1}^{t}{\mathsf{Pr}}\!\left[\begin{aligned} i_{0}\text{ was sampled $m-1$ times before round }j_{0}\\ \text{ and }\sigma_{0}\mbox{ is sampled in round $j_{0}$ from }S_{i_{0}}\end{aligned};{\mathsf{U}}_{1,\ldots,k}\right]=\sum\limits_{m=1}^{t}{\mathsf{Pr}}\!\left[S_{i_{0}}[m]=\sigma_{0};{{\mathsf{U}}_{1,\ldots,k}}\right]$
+$=$ $\sum\limits_{\omega}\sum\limits_{m=1}^{t}{\mathsf{Pr}}\!\left[\omega[m]=\sigma_{0}\mid S_{i_{0}}=\omega\right]\cdot{\mathsf{Pr}}\!\left[U_{i_{0}}=\omega\right]=\sum\limits_{\omega}{\mathsf{Pr}}\!\left[U_{i_{0}}=\omega\text{ and }\omega[m]=\sigma_{0}\right]=\frac{1}{|T_{i_{0}}|}$
+Therefore, $\frac{1}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq{\mathsf{Pr}}\!\left[(\sigma_{0},i_{0})\mbox{ is sampled in round }j_{0};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq(1+\epsilon_{\textsf{sz}})\frac{1}{\sum_{i=1}^{k}\textsf{sz}_{i}}.$
+For any $i_{0}\in[t]$ and $\sigma_{0}\in S_{i_{0}}$, we have ${\mathsf{Pr}}\!\left[(\sigma_{0},i_{0})\mbox{ is sampled };{{\mathsf{U}}_{1,\ldots,k}}\right]$
+[…]
+$=$ $\frac{\textsf{sz}_{i_{0}}}{\sum_{i}^{k}\textsf{sz}_{i}}\times\frac{1}{|T_{i_{0}}|}$
+Probability that $i_{0}$ is sampled in Line 1 is independent of the distributions of the samples.
+The last equality is because we have assumed $\forall i$, $|S_{i}|\geq t$, so the if condition in Line 1 is always satisfied.
+Hence once $i_{0}$ is sampled in Line 1 probability that $\sigma_{0}$ is obtained is exactly $1/|T_{i_{0}}|$, since the set $S_{i_{0}}$ contains elements sampled uniformly (with replacement) from $T_{i_{0}}$.
+Therefore, $\frac{1}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq{\mathsf{Pr}}\!\left[\sigma_{0},i_{0})\mbox{ is sampled };{{\mathsf{U}}_{1,\ldots,k}}\right]\leq(1+\epsilon_{\textsf{sz}})\frac{1}{\sum_{i=1}^{k}\textsf{sz}_{i}}.$
+Let $Y_{j}$ be the random variable denoting whether the counter $Y$ is increased in the $j$th iteration of the loop.
+Note that $Y_{j}=1$ if the pair $(\sigma_{0},i_{0})$ sampled in the $j$th iteration is in $\mathcal{U}_{\textsf{unique}}$.
+This is because in Line 1 it is checked if the sampled pair is in $\mathcal{U}_{\textsf{unique}}$.
+So, $\mathbb{E}[Y_{j};{\mathsf{U}}_{1,\ldots,k}]=\sum_{(\sigma,i)\in\mathcal{U}_{\textsf{unique}}}{\mathsf{Pr}}\!\left[(\sigma,i)\mbox{ is sampled in round }j;{{\mathsf{U}}_{1,\ldots,k}}\right]$; here the ‘;’ in the expectation has an analogous definition to the case when used for probability.
+Thus,
+$\frac{|\mathcal{U}_{\textsf{unique}}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\mathbb{E}[Y_{j};{\mathsf{U}}_{1,\ldots,k}]\leq(1+\epsilon_{\textsf{sz}})\frac{|\mathcal{U}_{\textsf{unique}}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$
+Note that $Y=\sum_{i}^{t}Y_{i}$.
+Thus at the end of the algorithm the expected value of $Y$ is between $\frac{t|\mathcal{U}_{\textsf{unique}}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}$ and $(1+\epsilon_{\textsf{sz}})^{2}\frac{t|\mathcal{U}_{\textsf{unique}}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$.
+Now by Chernoff bound we have,
+${\mathsf{Pr}}\!\left[Y\not\in(1\pm\frac{\epsilon}{2})\mathbb{E}[Y_{j};{\mathsf{U}}_{1,\ldots,k}];{{\mathsf{U}}_{1,\ldots,k}}\right]\leq 2exp(-\frac{\epsilon^{2}\mathbb{E}[Y_{j};{\mathsf{U}}_{1,\ldots,k}]}{12}).$
+We also know that,
+$\mathbb{E}[Y;{\mathsf{U}}_{1,\ldots,k}]\geq$ $\frac{t|\mathcal{U}_{\textsf{unique}}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}|T_{i}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}\textsf{sz}_{i}}{(1+\epsilon_{\textsf{sz}})^{2}\sum_{i=1}^{k}\textsf{sz}_{i}}\geq 12\log\left(\frac{4}{\delta}\right)$
+The last inequality is due to the setting of the parameter $t$.
+Thus, by Chernoff bound, probability that the output of the algorithm is not between $(1-\frac{\epsilon}{2})\frac{|\mathcal{U}_{\textsf{unique}}|}{(1+\epsilon_{\textsf{sz}})}$ and $(1+\frac{\epsilon}{2})(1+\epsilon_{\textsf{sz}})|\mathcal{U}_{\textsf{unique}}|$ is at most $\delta/2$.
+Since $(1-\frac{\epsilon}{2})\geq\frac{1}{1+\epsilon}$, so we have ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq\frac{\delta}{2}$.
 […]
 […]
-$=$ $\frac{\textsf{sz}_{i_{0}}}{\sum_{i}^{k}\textsf{sz}_{i}}\times\frac{1}{|T_{i_{0}}|}$
-The last equality is because we have assumed $\forall i$, $|S_{i}|\geq t$, so the if condition in line 2 is always satisfied.
-Hence once $i_{0}$ is sampled in line 2 probability that $\sigma_{0}$ is obtained is exactly $1/|T_{i_{0}}|$, since the set $S_{i_{0}}$ contains elements sampled uniformly (with replacement) from $T_{i_{0}}$.
-Therefore,
-$\frac{1}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\Pr[(\sigma_{0},i_{0})\mbox{ is sampled }]\leq(1+\epsilon_{\textsf{sz}})\frac{1}{\sum_{i=1}^{k}\textsf{sz}_{i}}.$
-Let $Y_{j}$ be the random variable denoting whether the counter $Y$ is increased in the $j$th iteration of the loop.
-Note that $Y_{j}=1$ if the pair $(\sigma_{0},i_{0})$ sampled in the $j$th iteration is in $\mathcal{U}_{1}$.
-This is because in line 2 it is checked if the sampled pair is in $\mathcal{U}_{1}$.
-So, $\mathbb{E}[Y_{j}]=\sum_{(\sigma,i)\in\mathcal{U}_{1}}\Pr[(\sigma,i)\mbox{ is sampled}]$. Thus,
-$\frac{|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\mathbb{E}[Y_{j}]\leq(1+\epsilon_{\textsf{sz}})\frac{|\mathcal{U}_{1}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$
-Note that $Y=\sum_{i}^{t}Y_{i}$.
-Thus at the end of the algorithm the expected value of $Y$ is between $\frac{t|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}$ and $(1+\epsilon_{\textsf{sz}})^{2}\frac{t|\mathcal{U}_{1}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$.
-Now by Chernoff bound we have
-$\Pr[Y\not\in(1\pm\frac{\epsilon}{2})\mathbb{E}[Y]]\leq 2exp(-\frac{\epsilon^{2}\mathbb{E}[Y]}{12}).$
-We also know that,
-$\mathbb{E}[Y]\geq$ $\frac{t|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}|T_{i}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}\textsf{sz}_{i}}{(1+\epsilon_{\textsf{sz}})^{2}\sum_{i=1}^{k}\textsf{sz}_{i}}$
-$\geq$ $12\log(\frac{4}{\delta})$
-The last inequality is due to the setting of the parameter $t$.
-Thus, by Chernoff bound, probability that the output of the algorithm is not between $(1-\frac{\epsilon}{2})\frac{|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})}$ and $(1+\frac{\epsilon}{2})(1+\epsilon_{\textsf{sz}})|\mathcal{U}_{1}|$ is at most $\delta/2$.
-Since $(1-\frac{\epsilon}{2})\geq\frac{1}{1+\epsilon}$, so we have $\Pr[\mathsf{Fail}\ \cap\ \overline{\mathsf{Bad}}]\leq\frac{\delta}{2}$.
-Part 2: If for all $i$, $|S_{i}|\geq\texttt{thresh}$
-Note that, the else clause in line 2 to 2 is ever reached, if for some $i$, the number of times $i$ is sampled in the $t$ iterations is more than thresh.
-Let $\mathsf{Bad}_{i}$ denote the event that the index $i$ is sampled (in line 2) more that thresh number of times during the course of the algorithm, that is during the $t$ runs of the loop (lines 2-2).
+Let $\mathsf{Bad}_{i}$ denote the event that the index $i$ is sampled (in Line 1) more that thresh number of times during the course of the algorithm, that is during the $t$ runs of the loop (Lines 1-1).
 And let the event $\mathsf{Bad}$ be $\cup_{i=1}^{k}\mathsf{Bad}_{i}$.
 We will now upper bound, $\Pr[\mathsf{Bad}_{i}]$.
-Let $W_{i}$ be the random variable counting the number times $i$ is sampled.
-$\mathbb{E}[W_{i}]=t\times\frac{\textsf{sz}_{i}}{\sum_{i=1}^{k}\textsf{sz}_{i}}$ which is less than $\frac{12(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\log(4/\delta)$.
+Let $W_{i}$ be the random variable counting the number times $i$ is sampled: $\mathbb{E}[W_{i}]=t\times\frac{\textsf{sz}_{i}}{\sum_{i=1}^{k}\textsf{sz}_{i}}$ which is less than $\frac{12(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\log(4/\delta)$.
 By simple application of Chernoff Bound we see that
-$\Pr[W_{i}\geq(1+\Delta)\mathbb{E}[W_{i}]]\leq exp(-\frac{\Delta^{2}\mathbb{E}[W_{i}]}{2+\Delta})\leq exp(-\frac{\Delta\mathbb{E}[W_{i}]}{2}),$
+$\Pr[W_{i}\geq(1+\Delta)\mathbb{E}[W_{i}]]\leq exp\left(-\frac{\Delta^{2}\mathbb{E}[W_{i}]}{2+\Delta}\right)\leq exp\left(-\frac{\Delta\mathbb{E}[W_{i}]}{2}\right),$
 the last inequality is for $\Delta\geq 2$.
@@
 So by union bound $\Pr[\mathsf{Bad}]\leq\delta/2$.
-5. Correctness of Algorithm 3
-In the following three sections, we prove the three parts of Theorem 2 individually.
-5.1. Proof of Theorem 2(1)
-Consider an execution of Algorithm 3 and let $w=w_{1}w_{2}\ldots w_{\ell}\in\mathcal{L}({q}^{\ell})$ be the string constructed right before entering the branch at Line 3.
+Observe that for large $k$, $\Delta\geq 2$ as desired.
+$\Pr[\mathsf{Bad_{i}}]\leq\frac{\delta}{2k}$.Observe that for large $k$, $\Delta\geq 2$ as desired.
+So by union bound, we have: ${\mathsf{Pr}}\!\left[\mathsf{Bad};{{\mathsf{U}}_{1,\ldots,k}}\right]=\Pr[\mathsf{Bad}]\leq\delta/2.$
+4.2. Correctness of Algorithm 2
+We prove the three parts of Theorem 2 individually.
+Proof of Theorem 2(1)
+Consider an execution of Algorithm 2 and let $w=w_{1}w_{2}\ldots w_{\ell}\in\mathcal{L}({q}^{\ell})$ be the string constructed right before entering the branch at Line 2.
 Let $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}$ be the value of the variable $\mathsf{pr_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}$ at level $i$ and let $v$ be the value of $\varphi$ before the function returns.
@@
 $p_{w_{i},i}=\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}$
-Thus,
-(2) $v=\frac{\gamma_{0}}{\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}}=\gamma_{0}\cdot\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}{\textsf{sz}^{i}_{w_{i}}}$
+Thus, $v=\frac{\gamma_{0}}{\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}}=\gamma_{0}\cdot\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}{\textsf{sz}^{i}_{w_{i}}}$
 Recall that $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$.
-Observe that $\mathcal{L}(P^{i})=\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})\uplus\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})$, and thus,
-$|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})|+|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})|=|\mathcal{L}(P^{i})|$
+Observe that $\mathcal{L}(P^{i})=\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})\uplus\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})$, and thus, $|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})|+|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})|=|\mathcal{L}(P^{i})|$
 Hence, under the event $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}\cap\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ we have
@@
 $\begin{array}[]{c}v\leq\frac{2}{3e}{\cdot}(1+\beta)^{\ell(\ell+2)}\leq\frac{2}{3e}\cdot e\leq 1\\ \end{array}$
-It then follows that
-${\mathsf{Pr}}\!\left[\mathtt{Fail}_{1}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=0$
-Substituting, $v=\frac{\gamma_{0}}{p_{w_{\ell},\ell}\cdot p_{w_{\ell-1},\ell-1}\cdots p_{w_{1},1}}$, we then have
-$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]\\ =v\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\frac{\gamma_{0}}{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},\ell}\cdot p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},\ell-1}\cdots p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},1}}\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\gamma_{0}\end{array}$
-5.2. Proof of Theorem 2(2)
+It then follows that ${\mathsf{Pr}}\!\left[\mathtt{Fail}_{1}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]$ is equal to 00..
+Substituting, $v=\frac{\gamma_{0}}{p_{w_{\ell},\ell}\cdot p_{w_{\ell-1},\ell-1}\cdots p_{w_{1},1}}=\frac{\gamma_{0}}{\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}}$, we then have
+${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=v\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\frac{\gamma_{0}}{\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}}\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\gamma_{0}$
+Proof of Theorem 2(2)
 We will now estimate the (conditional) probability of the event $\mathtt{Fail}$.
@@
 $=1{-}\!\!\!\!\!\!\sum\limits_{w\in\mathcal{L}({q}^{\ell})}\!\!\!{\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]$
-$=1{-}\sum\limits_{w\in\mathcal{L}({q}^{\ell})}\gamma_{0}=1{-}\gamma_{0}|\mathcal{L}({q}^{\ell})|\leq 1-\frac{2}{3e}\cdot\frac{|\mathcal{L}({q}^{\ell})|}{\mathsf{N}({q}^{\ell})}$
-$\leq 1-\frac{2}{3e}\cdot\frac{1}{(1+\beta)^{\ell}}\leq 1-\frac{2}{3e^{2}}$
-5.3. Proof of Theorem 2(3)
+$=1{-}\!\!\!\!\!\!\sum\limits_{w\in\mathcal{L}({q}^{\ell})}\!\!\!\gamma_{0}=1-\gamma_{0}|\mathcal{L}({q}^{\ell})|\leq 1-\frac{2}{3e}\cdot\frac{|\mathcal{L}({q}^{\ell})|}{\mathsf{N}({q}^{\ell})}\leq 1-\frac{2}{3e}\cdot\frac{1}{(1+\beta)^{\ell}}\leq 1-\frac{2}{3e^{2}}$
+Proof of Theorem 2(3)
 We would like to use Theorem 1 to prove Theorem 2(3). Towards this, we first establish that the pre-conditions of Theorem 1 hold.
@@
 Thus for each ${q}^{i}$, we have
-    $\texttt{thresh}\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\leq|\mathsf{S}({q}^{i-1})|$
+    $\texttt{thresh}\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\leq|\mathsf{S}({q}^{i-1})|.$
   2. (2)
@@
   3. (3)
-    The event $\texttt{AccurateN}_{\leq\ell-1}$ demands that for each $i\leq\ell-1$, we have $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|$.
+    The event $\texttt{AccurateN}_{\leq\ell-1}$ demands that for each $i\leq\ell-1$, we have
+    $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|.$
 From Theorem 1, it thus follows that, for all $i$,
@@
 Therefore, applying union-bound, we have
-$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\\ \geq 1-2n\cdot\eta^{\prime}=1-\frac{\eta}{2}\end{array}$
-6. Correctness of Algorithm 4
-We first establish some preliminaries relevant to our proof.
-6.1. Total Variation Distance
-The total variation distance is a statistical distance between two discrete valued random variables $X,Y$ over domain $\Omega$ and is defined as follows:
-$\mathsf{TV}(X,Y)$ $=$ $\frac{1}{2}\sum_{v\in\Omega}\left|\Pr[X=v]-\Pr[Y=v]\right|$
-$=$ $\sum_{v\,:\,\Pr[X=v]\geq\Pr[Y=v]}\left(\Pr[X=v]-\Pr[Y=v]\right)\;.$
-We will use the following folklore result (c.f.
-Lemma 6.3 of  (Vadhan, 2012)).
-Here, for a random variable $X$ over the sample space $\Omega$, we use the notation ${\mathsf{Pr}}\!\left[E:X\right]$ to denote the probability of the event $E\subseteq\Omega$.
-Lemma 6.0.
-Let $X$ and $Y$ be the random variables taking values over a measurable set $\Omega$.
-Let $f:\Omega\to\hat{\Omega}$ be a (possibly randomized) function with measurable co-domain $\hat{\Omega}$.
-Let $E\subseteq\hat{\Omega}$ be an arbitrary event.
-Then, the following holds:
-$|{\mathsf{Pr}}\!\left[E:{f(X)}\right]-{\mathsf{Pr}}\!\left[E:f(Y)\right]|\leq\mathsf{TV}(X,Y)$
-6.2. Proof of Theorem 3
+$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-2n\cdot\eta^{\prime}=1-\frac{\eta}{2}\end{array}$
+4.3. Correctness of Algorithm 3
 We prove the main result (Theorem 3) by induction on the level $\ell$.
@@
 After this, we characterize the real computed estimates and samples by arguing that the distance of the corresponding random variables, from those computed using perfectly uniform samples, is small.
-We will be using the total variation distance and Lemma 1 to structure the proof.
 We first establish some helpful auxiliary results (Lemma 1 and Lemma 2) towards our main proof.
-As before, we will often invoke the the probability space obtained by replacing the sets $\mathsf{S}({q}^{j})$ by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$.
-In addition to the previously introduced notation ${\mathsf{Pr}}\!\left[;{{\mathsf{U}}_{\leq i}}\right]$, we will also use the notation ${\mathsf{TV}}(X,Y\;;{\mathsf{U}}_{\leq i})$ (resp.
-${\mathsf{TV}}(X,Y\;\;\big{|}\;\;\mathcal{F}\;;{\mathsf{U}}_{\leq i})$) to denote the total variational distance between the distribution induced by random variables $X,Y$ (resp.
-distribution induced by $X,Y$ conditioned on the event $\mathcal{F}$) when the $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 is replaced by a set $\mathsf{U}({q}^{j})$ obtained by unirofmly sampling (with replacement) $|\mathsf{S}({q}^{j})|$ many elements from $\mathcal{L}({q}^{j})$, for all $q\in Q$ and for all $1\leq j\leq i$.
-Lemma 6.0.
+As before, we will often invoke distribution entanglement of ${\sf S}_{\leq\ell-1}$ by ${\mathsf{U}}_{\leq\ell-1}$ for each level $\ell$.
+Lemma 1 ().
 For each $\ell\leq n$ and $q\in Q$, we have
@@
 Proof.
-[…]
-thresh $=24\cdot\frac{(1+\beta)^{2\ell-2}}{\beta^{2}}\cdot\log\left(\frac{4k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}{\eta^{\prime}}\right)\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{8m^{2}n}{\delta}\right)$
-$\leq\frac{384\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{8m^{2}}{\delta}\right)\leq\mathtt{ns}$
+Let us denote by $\mathcal{L}(\textsf{Pred}({q}^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))$ the set $\bigcup\limits_{p\in\textsf{Pred}({q}^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})}\mathcal{L}(p).$ Observe that,
+$\mathcal{L}({q}^{\ell})=\mathcal{L}(\textsf{Pred}({q}^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}))\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}\}\uplus\mathcal{L}(\textsf{Pred}({q}^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}))\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}.$
+Now, we would like to apply Theorem 1 to show that $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ (at level $\ell$) is a good estimate.
+To this end, we first show that the pre-conditions stated in Theorem 1 hold.
+This is because, (1) under the event $\texttt{AccurateN}_{\leq\ell-1}$, the estimates $\{\mathsf{N}({p}^{\ell-1})\}_{p\in Q}$ satisfy desired requirements.
+Next, (2) for each $i\leq\ell-1$, the samples $\{\mathsf{S}({p}^{\ell-1})\}_{p\in Q}$ also obey the desired requirements.
+This is because thresh (in Algorithm 1), after substituting $\beta,\eta$, satisfies
+thresh $=24\cdot\frac{(1+\beta)^{2\ell-2}}{\beta^{2}}\cdot\log\left(\frac{4k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}{\eta^{\prime}}\right)\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{8m^{2}n}{\delta}\right)\leq\frac{384\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{8m^{2}}{\delta}\right)\leq\mathtt{ns}$
 Therefore, from Theorem 1, we have
-${\mathsf{Pr}}\!\left[\begin{array}[]{l}\frac{|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|}{(1+\beta)^{\ell}}\leq\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\\ \quad\quad\leq(1+\beta)^{\ell}\;|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|\end{array}\big{|}\;\;\texttt{AccurateN}_{\ell-1};{{\mathsf{U}}_{\ell-1}}\right]\geq 1-\frac{\eta}{2}$
+${\mathsf{Pr}}\!\left[\frac{|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|}{(1+\beta)^{\ell}}\leq\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq(1+\beta)^{\ell}\;|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|\;\big{|}\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
+is at least $(1-\frac{\eta}{2})$.
 Since $\mathsf{N}({q}^{\ell})=\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$, we have
-${\mathsf{Pr}}\!\left[\begin{array}[]{l}\frac{|\mathcal{L}({q}^{\ell})|}{(1+\beta)^{\ell}}\leq\mathsf{N}({q}^{\ell})\\ \quad\quad\leq(1+\beta)^{\ell}|\mathcal{L}({q}^{\ell})|\end{array}\mid\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta$
-Lemma 6.0.
+${\mathsf{Pr}}\!\left[\frac{|\mathcal{L}({q}^{\ell})|}{(1+\beta)^{\ell}}\leq\mathsf{N}({q}^{\ell})\leq(1+\beta)^{\ell}|\mathcal{L}({q}^{\ell})|\;\big{|}\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta\qquad\qed$
+In addition to the previously introduced notation ${\mathsf{Pr}}\!\left[;{{\mathsf{U}}_{\leq i}}\right]$, we will also use the notation ${\mathsf{TV}}(X,Y\;;{\mathsf{U}}_{\leq i})$ (resp.
+${\mathsf{TV}}(X,Y\,\big{|}\,\mathcal{F}\;;{\mathsf{U}}_{\leq i})$) to denote the total variational distance between the distribution induced by random variables $X,Y$ (resp.
+distribution induced by $X,Y$ conditioned on the event $\mathcal{F}$) when ${\sf S}_{\leq i}$ is distribution entangled to ${\mathsf{U}}_{\leq i}$.
+Formally,
+${\mathsf{TV}}(X,Y\mid\mathcal{F}\;;{\mathsf{U}}_{\leq i})=\sum_{\omega^{\prime}}$ ${\mathsf{Pr}}\!\left[{\mathsf{U}}_{\leq i}=\omega^{\prime}\right]\left(\sum_{\omega}\Big{(}\Pr[X=\omega\mid\mathcal{F}\cap({\sf S}_{\leq i}=\omega^{\prime})]\right.$
+$-\min\left(\Pr[X=\omega\mid\mathcal{F}\cap({\sf S}_{\leq i}=\omega^{\prime})],\Pr[Y=\omega\mid\mathcal{F}\cap({\sf S}_{\leq i}=\omega^{\prime})]\right)\Big{)}\Big{)}$
+In the above expression, $\omega$ ranges over the union of domains of $X$ and $Y$ while $\omega^{\prime}$ ranges over the union of domains of ${\sf S}_{\leq i}$ and ${\mathsf{U}}_{\leq i}$.
+Lemma 2 ().
 For each $\ell\leq n$ and $q\in Q$, we have
-${\mathsf{TV}}(S(q^{\ell}),U(q^{\ell})\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\ell-1})\leq\eta$
+${\mathsf{TV}}(\mathsf{S}({q}^{\ell}),\mathsf{U}({q}^{\ell}))\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})\leq\eta$
 Proof.
 Fix $\ell$ and $q$.
-Let $\mathtt{Fail}^{j}$ denote the event that the $j^{\text{th}}$ call to sample (in the loop starting at Line 4) for $i=j$ returns $\bot$.
+Let $\mathtt{Fail}^{j}$ denote the event that the $j^{\text{th}}$ call to sample (in the loop starting at Line 3) for $i=j$ returns $\bot$.
 Similarly, let $\texttt{Accurate}\texttt{AppUnion}^{j}$ denote the event AccurateAppUnion corresponding to the call to sample for $i=j$.
-Using Iversion bracket notation, we denote the indicator variable $[\mathtt{Fail}^{j}]$.
+Using Iverson bracket notation (Iverson, 1962), we denote the indicator variable $[\mathtt{Fail}^{j}]$.
 Also, we use SmallS to denote the event that $\alpha_{{q}^{\ell}}>0$.
 Observe that SmallS occurs if and only if $\sum\limits_{j=1}^{\mathtt{xns}}[\mathtt{Fail}^{j}]\geq\mathtt{xns}-\mathtt{ns}$.
-Recall, we have, for each $j$
-${\mathsf{Pr}}\!\left[[\mathtt{Fail}^{j}]\;\;\big{|}\;\;\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq 1-\frac{2}{3e^{2}}$
+Recall, we have, for each $j$, ${\mathsf{Pr}}\!\left[[\mathtt{Fail}^{j}]\;\;\big{|}\;\;\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$ is at most $(1-\frac{2}{3e^{2}})$.
 Therefore, from Chernoff bound, we have
-${\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\;\;\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\frac{\eta}{2}$
+${\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\;\;\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\frac{\eta}{2}.$
 Furthermore, from Theorem 2 (part 3), we have
-${\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\eta/2$
-Note that assuming $\texttt{AccurateN}_{\leq\ell}$ and ${\mathsf{U}}_{\leq\ell-1}$, is the event SmallS does not happen and for all $j$ $\texttt{Accurate}\texttt{AppUnion}^{j}$ holds then the distribution of $S(q^{\ell})$ and $U(q^{\ell})$ is identical.
+${\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\frac{\eta}{2}.$
+Note that given the event $\texttt{AccurateN}_{\leq\ell}$ and under the distributions obtained using ${\mathsf{U}}_{\leq\ell-1}$, if the event SmallS does not happen and the event $\bigcap_{j}\texttt{Accurate}\texttt{AppUnion}^{j}$ holds then the distribution of $\mathsf{S}(q^{\ell})$ and $\mathsf{U}(q^{\ell})$ is identical.
 So,
@@
 $\leq{\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
-$\quad+{\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
-$\leq\eta/2+\eta/2=\eta$
-Now, we are ready to state and prove the main theorem.
-Proof of Theorem 3.
-For ease of exposition, we will refer to the event $\texttt{AccurateN}_{n}$ in our analysis. Applying  Lemma 1, we have
-${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq$
-$\sum_{\ell\in[1,n];q\in Q}{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
-$\qquad+\sum\limits_{\ell\in[1,n-1];q\in Q}\mathsf{TV}(S(q^{\ell}),U(q^{\ell})\mid\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
-By Lemma 2 and Lemma 3 and applying union bound oven all $\ell\in[1,n]$ and all $q\in Q$ we have
+$\quad\qquad\qquad+{\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\eta/2+\eta/2=\eta\qed$
+4.3.1. Proof of Theorem 3
+Our proof of the Theorem 3 relies on upper bounding the probability with which the estimate of Algorithm 3 is outside the desired range.
+For this, we will set up new random variables and relate its joint distribution with the statsitical distance between the samples constructed by the algorithm and the ideal set of uniform samples.
+Recall that $\{\mathsf{S}({q}^{\ell})\}_{\ell\in[0,n],q\in Q}$ represent the set of random variables corresponding to the (multi)sets of samples we obtain during run of Algorithm 3.
+Furthermore, $\{\mathsf{U}({q}^{\ell})\}_{\ell\in[0,n],q\in Q}$ represent the set of random variables that correspond to repeatedly sampling $\mathtt{ns}$ elements (with replacement) uniformly at random from $\mathcal{L}({q}^{\ell})$.
+It is worth observing that while, for all $q,p,\ell,\ell^{\prime}$, $\mathsf{S}({q}^{\ell})$ and $\mathsf{S}({p}^{\ell^{\prime}})$ are dependent but $\mathsf{U}({q}^{\ell})$ and $\mathsf{U}({q}^{\ell^{\prime}})$ are independent.
+Also, the definition of distribution $\mathsf{U}({q}^{\ell})$ is independent of  Algorithm 3 as it is defined solely based on the set of $\mathcal{L}(\mathsf{S}({q}^{\ell}))$.
+Since we are talking about the sets of random variables, there is an implicit assumption that these sets are ordered.
+We will now define two sequences of random variables:
+$X=\{X^{q}_{\ell}\}_{\ell\in[0,n],q\in Q}\ \mbox{ and }\ Y=\{Y^{q}_{\ell}\}_{\ell\in[0,n],q\in Q},$
+Instead of explicitly specifying them, we will instead specify properties about their joint distribution.
+We will use the notation $X_{\leq i}$ (and resp.
+$Y_{\leq i})$) to denote the ordered set $\{X^{q}_{\ell}\}_{\ell\in[1,i],q\in Q}$ (resp.
+the ordered set $\{Y^{q}_{\ell}\}_{\ell\in[1,i],q\in Q}$).
+The joint distribution of $(X,Y)$ has the following properties:
+  1. (1)
+    For every $q\in Q$, $X^{q}_{0}=Y^{q}_{0}={\sf S}^{q}_{0}$.
+  2. (2)
+    For every $q\in Q,\ell\in[1,n]$, $X^{q}_{\ell}=S(q^{\ell})$.
+  3. (3)
+    For all sets $\omega$, sequences of sets $\omega^{\prime}$, $\ell\in[1,n]$, and $q\in Q$, we have
+    ${\mathsf{Pr}}\!\left[X^{q}_{\ell}=Y^{q}_{\ell}=\omega\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]$
+    $\qquad\qquad=\min\left({\mathsf{Pr}}\!\left[U(q^{\ell})=\omega\right],{\mathsf{Pr}}\!\left[X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right)$
+    ${\mathsf{Pr}}\!\left[X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\leq{\mathsf{Pr}}\!\left[X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]$
+We remark that such a joint distribution always exists.
+This is because (3) only restricts the case when $X_{\ell}^{q}=Y_{\ell}^{q}$, thereby, providing enough choices for assignments of probability values to the cases where $X_{\ell}^{q}\neq Y_{\ell}^{q}$ such that $Y^{q}_{\ell}$ has a well-defined probability distribution.
+In Appendix A.1 we give a concrete distribution that realizes these properties.
+We will begin with the following simple observation, whose proof is deferred to Appendix.
+Claim 3 ().
+${\mathsf{Pr}}\!\left[(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\cap\texttt{AccurateN}_{\leq\ell-1}\right]\leq{\mathsf{Pr}}\!\left[{\mathsf{U}}_{\leq\ell-1}=\omega\right]$
+Next, we relate the random variables defined above to the desired statistical distance:
+Claim 4 ().
+${\mathsf{Pr}}\!\left[(X_{\ell}^{q}\neq Y_{\ell}^{q})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap\texttt{AccurateN}_{\leq\ell}\right]$ $\leq{\mathsf{TV}}(\mathsf{S}({q}^{\ell}),\mathsf{U}({q}^{\ell})\;\big{|}\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
+Proof.
+${\mathsf{Pr}}\!\left[(X_{\ell}^{q}\neq Y_{\ell}^{q})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap\texttt{AccurateN}_{\leq\ell}\right]$
+$=\sum_{\omega,\omega^{\prime}}{\mathsf{Pr}}\!\left[(X_{\ell}^{q}=\omega)\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]$
+$\quad-\sum_{\omega,\omega^{\prime}}{\mathsf{Pr}}\!\left[(X_{\ell}^{q}=Y_{\ell}^{q}=\omega)\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]$
+$=\sum_{\omega^{\prime}}{\mathsf{Pr}}\!\left[(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\cdot$
+$\qquad\left(\sum_{\omega}{\mathsf{Pr}}\!\left[(X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right.$
+$\left.\qquad-\sum_{\omega}{\mathsf{Pr}}\!\left[(X_{\ell}^{q}=Y_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right)$
+$\leq\sum_{\omega^{\prime}}{\mathsf{Pr}}\!\left[U_{\leq\ell-1}=\omega^{\prime}\right]\cdot\big{(}\sum_{\omega}\left({\mathsf{Pr}}\!\left[(X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right.$
+$\qquad\qquad-\min\left({\mathsf{Pr}}\!\left[U(q^{\ell})=\omega\right],\left.\left.{\mathsf{Pr}}\!\left[X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right)\right)\right)$
+$={\mathsf{TV}}(\mathsf{S}({q}^{\ell}),\mathsf{U}({q}^{\ell})\;\big{|}\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
+We now focus on deriving an upper bound on $\texttt{AccurateN}_{n}$ as stated in the following claim, whose proof is deferred to Appendix.
+Claim 5 ().
+$\overline{\texttt{AccurateN}_{n}}\subseteq\bigcup_{\ell=1}^{n-1}(\overline{\texttt{AccurateN}_{\ell}}\cap{\texttt{AccurateN}_{\leq\ell-1}}\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}))$
+$\qquad\qquad\qquad\qquad\cup\bigcup_{\ell=1}^{n-1}\left((X_{\ell}\neq Y_{\ell})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell}}\right)$
+We can now finish the proof of Theorem 3. We first observe that:
+${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{\ell}}\cap{\texttt{AccurateN}_{\leq\ell-1}}\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\right]$
+$\leq\sum_{\omega,q}{\mathsf{Pr}}\!\left[\texttt{AccurateN}_{\leq\ell-1}\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\right]\times$
+$\qquad\qquad\qquad\qquad{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\cap{\texttt{AccurateN}_{\leq\ell-1}}\right]$
+$\leq\sum_{\omega,q}{\mathsf{Pr}}\!\left[U_{\leq\ell-1}=\omega\right]{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\cap{\texttt{AccurateN}_{\leq\ell-1}}\right]$
+$=\sum_{\omega,q}{\mathsf{Pr}}\!\left[\texttt{AccurateN}_{\leq\ell-1}=\omega\right]{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\mid({\sf S}_{\leq\ell-1}=\omega\cap Y_{\leq\ell-1}=\omega\cap{\texttt{AccurateN}_{\leq\ell-1}}\right]$
+$\leq\sum_{q}{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
+The final step in the derivation uses the observation that given ${\sf S}_{\leq\ell-1}$ and ${\texttt{AccurateN}_{\leq\ell-1}}$, the probability of the event $\texttt{AccurateN}_{q,\ell}$ is only dependent on the randomness of the algorithm AppUnion.
+Thus, given ${\sf S}_{\leq\ell-1}$ and ${\texttt{AccurateN}_{\leq\ell-1}}$, the event $\texttt{AccurateN}_{q,\ell}$ is conditionally independent of $Y_{\leq\ell-1}$.
+Putting together, we have:
+${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq$ $\sum_{\ell\in[1,n];q\in Q}{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
+$\qquad+\sum\limits_{\ell\in[1,n-1];q\in Q}{\mathsf{TV}}(\mathsf{S}({q}^{\ell}),\mathsf{U}({q}^{\ell})\mid\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
+By Lemma 1 and Lemma 2 and applying union bound over all $\ell\in[1,n]$ and all $q\in Q$ we have
 ${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq n\cdot m\cdot\eta+n\cdot m\cdot\eta\leq\delta.$
@@
 Now for each $\ell$ and $q$ there is one call to $\texttt{AppUnion}_{\beta,\eta}$ with the number of sets being $O(m)$ and $\mathtt{xns}$ calls to sample().
-So by Theorem 1 the time taken by AppUnion is $O(m/\beta^{2}\log(m/\eta))$ number of calls to the oracle.
-So the total complexity is $\widetilde{O}(mn^{4}/\epsilon^{2}\log(1/\delta))$, ignoring the factors of $\log(n+m)$.
-The sample() on the other hand is a recursive function which calls $\texttt{AppUnion}_{\beta,\eta}()$ every recursion and the recursion happens $O(n)$ times.
-So the time complexity for each of the $\mathtt{xns}$ calls to sample() is $\widetilde{O}(mn^{5}/\epsilon^{2}\log(1/\delta))$ calls to the oracle.
-Now note that the membership oracle takes $O(m^{2}n)$ steps for each call.
-So in total the time complexity of the Algorithm 4 is $\widetilde{O}(mn\cdot\mathtt{xns}\cdot\frac{mn^{5}}{\epsilon^{2}}\log(1/\delta)\cdot(m^{2}n))=\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.
+So by Theorem 1, the time taken by AppUnion is the time it takes to do $O(m/\beta^{2}\log(m/\eta))$ number of calls to the membership oracle.
+So the number of calls to the membership oracle is $\widetilde{O}(mn^{4}/\epsilon^{2}\log(1/\delta))$, ignoring the factors of $\log(n+m)$.
+The method sample() on the other hand is a recursive function which calls $\texttt{AppUnion}_{\beta,\eta}()$ in every recursive step and the depth of the recursion is $O(n)$.
+So the time complexity for each of the $\mathtt{xns}$ calls to sample() is the time it takes for $\widetilde{O}(mn^{5}/\epsilon^{2}\log(1/\delta))$ calls to the oracle.
+Now, we note that the time complexity of the membership calls can be amortized as follows.
+First, for every string $w\in\bigcup\limits_{q,\ell}\{\mathsf{S}({q}^{\ell})\}$, we pre-compute and store the set of reachable states of $w$ in $\mathcal{A}$ in time $m^{2}|w|\in O(m^{2}n)$ time.
+This takes a total of $\widetilde{O}(m^{2}n\cdot mn^{5}/\epsilon^{2}\log(1/\delta))=\widetilde{O}(m^{3}n^{6}/\epsilon^{2}\log(1/\delta))$ time.
+The membership checks (in Algorithm 1, Line 1) can now be performed in $O(1)$ time.
+So in total the time complexity of the Algorithm 3 is $\widetilde{O}(mn\cdot\mathtt{xns}\cdot\frac{mn^{5}}{\epsilon^{2}}\log(1/\delta)+m^{3}n^{6}/\epsilon^{4}\log(1/\delta))$ which is $\widetilde{O}((m^{2}n^{10}+m^{3}n^{6})\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.
 5. Conclusions and Future Work
 We consider the approximate counting problem #NFA and propose a fully polynomial time randomized approximation scheme (FPRAS) that significantly improves the prior FPRAS (Arenas et al., 2019).
-Our algorithm deeply exploits recent progress in the streaming algorithms literature.
-Given the wide range of applications of counting and sampling from the language of an NFA in program analysis, testing and more broadly in Computer Science, we envision that further improvements in the complexity of approximating #NFA is a worthwhile avenue for future work.
+Given the wide range of applications of counting and sampling from the language of an NFA in applications including databases, program analysis, testing and more broadly in Computer Science, we envision that further improvements in the complexity of approximating #NFA is a worthwhile avenue for future work.
 Acknowledgements.
-The authors thank Pierre Snellart and Antoine Amarilli for introducing us to the #NFA problem, and are grateful to N.V.
-Vinodchandran and Pavan Aduri for insightful discussions on the Karp-Luby method for the union of sets.
-This work was supported in part by National Research Foundation Singapore under its NRF Fellowship Programme[NRF-NRFFAI1-2019-0004 ], Ministry of Education Singapore Tier 2 grant MOE-T2EP20121-0011, and Ministry of Education Singapore Tier 1 grants [R-252-000-B59-114] and [R-252-000-C28-133].
+We want to express our sincerest gratitude to Weiming Feng for carefully reviewing the proofs and identifying several issues present in earlier versions of the paper.
+We are grateful to Yash Pote, Uddalok Sarkar, and the anonymous reviewers for their constructive feedback, which enhanced the presentation and brought attention to relevant prior research.
+Umang Mathur was partially supported by a Singapore Ministry of Education (MoE) Academic Research Fund (AcRF) Tier 1 grant.
+Kuldeep Meel’s work was partly supported in part by the National Research Foundation Singapore [NRF-NRFFAI1-2019-0004], Ministry of Education Singapore Tier 2 grant MOE-T2EP20121-0011, and Ministry of Education Singapore Tier 1 Grant [R-252-000-B59-114].
+Appendix A Details from Section 4.3
+A.1. Joint Distribution on $X$ and $Y$
+Here, we will outline a concrete joint distribution that satisfies the properties we describe in Section 4.3.
+We will present the definition inductively, first defining the slice $\{X^{q}_{1},Y^{q}_{1}\}_{q\in Q}$, and then use the slices $\{X^{q}_{j},Y^{q}_{j}\}_{q\in Q,0\leq j\leq i}$ to define the slice $\{X^{q}_{i+1},Y^{q}_{i+1}\}_{q\in Q}$.
+Our inductive definition will also allow us to establish the properties we outlined.
+In the following, we use the notation $\Omega_{X,i,q}$ (resp.
+$\Omega_{Y,i,q}$) to denote the domain of $X^{q}_{i}$ (resp.
+$Y^{q}_{i}$).
+Analogously, we will use the notations $\Omega_{X,i}$, $\Omega_{Y,i}$, $\Omega_{X,\leq i}$ and $\Omega_{Y,\leq i}$ to denote respectively $\Pi_{q\in Q}\Omega_{X,i,q}$, $\Pi_{q\in Q}\Omega_{Y,i,q}$, $\Pi_{j=1}^{i}\Omega_{X,j}$, and $\Pi_{j=1}^{i}\Omega_{Y,j}$.
+Base Case(i=0).:
+Follows from the definition.
+Inductive Case.:
+We will define the distribution inductively by defining the conditional probabilities ${\mathsf{Pr}}\!\left[X^{q}_{i+1}=\omega_{X,i+1,q}\cap Y^{q}_{i+1}=\omega_{Y,i+1,q}|X_{\leq i}=\omega_{X,\leq i}\cap Y_{\leq i}=\omega_{Y,\leq i}\cap\texttt{AccurateN}_{\leq i}\right]$ and
+${\mathsf{Pr}}\!\left[X^{q}_{i+1}=\omega_{X,i+1,q}\cap Y^{q}_{i+1}=\omega_{Y,i+1,q}|X_{\leq i}=\omega_{X,\leq i}\cap Y_{\leq i}=\omega_{Y,\leq i}\cap\overline{\texttt{AccurateN}_{\leq i}}\right]$.
+For each $A\in\{\texttt{AccurateN}_{\leq i},\overline{\texttt{AccurateN}_{\leq i}}\}$, we have:
+$\begin{array}[]{l}{\mathsf{Pr}}\!\left[X^{q}_{i+1}=\omega_{X,i+1,q}\cap Y^{q}_{i+1}=\omega_{Y,i+1,q}|X_{\leq i}=\omega_{X,\leq i}\cap Y_{\leq i}=\omega_{Y,\leq i}\cap A\right]\\ \quad\quad=\begin{cases}\min\left(\begin{array}[]{l}{\mathsf{Pr}}\!\left[\mathsf{U}({q}^{i+1})=\omega_{X,i+1,q}\right],\\ {\mathsf{Pr}}\!\left[\mathsf{S}({q}^{i+1})=\omega_{X,i+1,q}|{\sf S}_{\leq i}=\omega_{X,\leq i}\cap A\right]\end{array}\right)&\text{if }\omega_{X,i+1,q}=\omega_{Y,i+1,q}\\ \\ \begin{array}[]{c}\left[\begin{array}[]{l}{\mathsf{Pr}}\!\left[\mathsf{S}({q}^{i+1})=\omega_{X,i+1,q}|{\sf S}_{\leq i}=\omega_{X,\leq i}\cap A\right]\\ -\\ \min\left(\begin{array}[]{l}{\mathsf{Pr}}\!\left[\mathsf{U}({q}^{i+1})=\omega_{X,i+1,q}\right],\\ {\mathsf{Pr}}\!\left[\mathsf{S}({q}^{i+1})=\omega_{X,i+1,q}|{\sf S}_{\leq i}=\omega_{X,\leq i}\cap A\right]\end{array}\right)\end{array}\right]\\ \hline\cr|\Omega_{Y,i+1,q}|-1\end{array}&\text{otherwise}\end{cases}\end{array}$
+Let us first observe that for each $A\in\{\texttt{AccurateN}_{\leq i},\overline{\texttt{AccurateN}_{\leq i}}\}$ and $\omega_{X,\leq i}$, we have
+${\mathsf{Pr}}\!\left[X_{i+1}^{q}=\omega_{X,i+1,q}|{X}_{\leq i}=\omega_{X,\leq i}\cap A\right]$
+$=$ ${\mathsf{Pr}}\!\left[\mathsf{S}({q}^{i+1})=\omega_{X,i+1,q}|{\sf S}_{\leq i}=\omega_{X,\leq i}\cap A\right]$
+Also, it is clear that $\sum_{\omega_{X,\leq i+1},\omega_{Y,\leq i+1}}{\mathsf{Pr}}\!\left[X_{\leq i+1}=\omega_{X,\leq i+1}\cap Y_{\leq i+1}=\omega_{Y,\leq i+1}\right]=1$.
+It is easy to see that the above joint distribution satisfies both the desired properties.
+A.2. Proof of Claim 3
+See 3
+Proof.
+${\mathsf{Pr}}\!\left[(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\cap\texttt{AccurateN}_{\leq\ell-1}\right]$
+$=$ ${\mathsf{Pr}}\!\left[(X_{1}=Y_{1}=\omega_{1})\cap\texttt{AccurateN}_{1}\right]\times$
+$\quad{\mathsf{Pr}}\!\left[(X_{2}=Y_{2}=\omega_{2})\cap\texttt{AccurateN}_{2}\mid(X_{1}=Y_{1}=\omega_{1})\cap\texttt{AccurateN}_{1}\right]\times$
+$\quad\quad{\mathsf{Pr}}\!\left[(X_{3}=Y_{3}=\omega_{3})\cap\texttt{AccurateN}_{3}\mid(X_{\leq 2}=Y_{\leq 2}=\omega_{1}\omega_{2})\cap\texttt{AccurateN}_{\leq 2}\right]\times\dots$
+$=$ $\prod_{k=1}^{\ell}{\mathsf{Pr}}\!\left[(X_{k}=Y_{k}=\omega_{k})\cap\texttt{AccurateN}_{k}\mid(X_{\leq(k-1)}=Y_{\leq(k-1)}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq(k-1)}\right]$
+Now, since $\Pr[A\cap B\mid C]=\frac{\Pr[A\cap B\cap C]}{\Pr[C]}\leq\frac{\Pr[A\cap B\cap C]}{\Pr[B\cap C]}=\Pr[A\mid B\cap C]$ so,
+${\mathsf{Pr}}\!\left[(X_{k}=Y_{k}=\omega_{k})\cap\texttt{AccurateN}_{k}\mid(X_{\leq(k-1)}=Y_{\leq(k-1)}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq(k-1)}\right]$
+$\leq$ ${\mathsf{Pr}}\!\left[(X_{k}=Y_{k}=\omega_{k})\mid(X_{\leq(k-1)}=Y_{\leq(k-1)}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq(k)}\right]$
+So,
+${\mathsf{Pr}}\!\left[(X_{k}=Y_{k}=\omega_{k})\cap\texttt{AccurateN}_{k}\mid(X_{\leq(k-1)}=Y_{\leq(k-1)}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq(k-1)}\right]$
+$\leq$ ${\mathsf{Pr}}\!\left[(X_{k}=Y_{k}=\omega_{k})\mid(X_{\leq(k-1)}=Y_{\leq(k-1)}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq(k)}\right]$
+From the second condition of the joint random variables $X$ and $Y$ we have
+${\mathsf{Pr}}\!\left[(X_{k}=Y_{k}=\omega_{k})\mid(X_{\leq(k-1)}=Y_{\leq(k-1)}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq(k)}\right]\leq$ ${\mathsf{Pr}}\!\left[U_{\leq k}=\omega\right]$
+So we have,
+${\mathsf{Pr}}\!\left[(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\cap\texttt{AccurateN}_{\leq\ell-1}\right]$
+$=$ $\prod_{k=1}^{\ell}{\mathsf{Pr}}\!\left[(X_{k}=Y_{k}=\omega_{k})\cap\texttt{AccurateN}_{k}\mid(X_{\leq(k-1)}=Y_{\leq(k-1)}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq(k-1)}\right]$
+$\leq$ $\prod_{k=1}^{\ell}{\mathsf{Pr}}\!\left[(X_{k}=Y_{k}=\omega_{k})\mid(X_{\leq(k-1)}=Y_{\leq(k-1)}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq(k)}\right]$
+$\leq$ $\prod_{k=1}^{\ell}{\mathsf{Pr}}\!\left[U_{\leq k}=\omega\right]={\mathsf{Pr}}\!\left[U_{\leq\ell-1}=\omega\right],$
+the last equality since the random variables $U_{k}$ are are uniformly generated and hence independent. ∎
+A.3. Proof of Claim 5
+See 5
+Proof.
+$\overline{\texttt{AccurateN}_{n}}\subseteq$ $(\overline{\texttt{AccurateN}_{n}}\cap(X_{\leq n-1}=Y_{\leq n-1}))\cup(X_{\leq n-1}\neq Y_{\leq n-1})$
+Expanding further, we have
+$\overline{\texttt{AccurateN}_{n}}\subseteq$ $(\overline{\texttt{AccurateN}_{n}}\cap{\texttt{AccurateN}_{\leq n-1}}\cap(X_{\leq n-1}=Y_{\leq n-1}))$
+$\cup(\overline{\texttt{AccurateN}_{n}}\cap\overline{\texttt{AccurateN}_{\leq n-1}}\cap(X_{\leq n-1}=Y_{\leq n-1}))$
+$\cup((X_{\leq n-1}\neq Y_{\leq n-1})\cap\overline{\texttt{AccurateN}_{\leq n-1}})\cup((X_{\leq n-1}\neq Y_{\leq n-1})\cap{\texttt{AccurateN}_{\leq n-1}})$
+Observing $(\overline{\texttt{AccurateN}_{n}}\cap\overline{\texttt{AccurateN}_{\leq n-1}})\subseteq\overline{\texttt{AccurateN}_{\leq n-1}}$ and then combining second and third terms, we have
+$\overline{\texttt{AccurateN}_{n}}\subseteq$ $(\overline{\texttt{AccurateN}_{n}}\cap{\texttt{AccurateN}_{\leq n-1}}\cap(X_{\leq n-1}=Y_{\leq n-1}))\cup\overline{\texttt{AccurateN}_{\leq n-1}}$
+$\cup((X_{\leq n-1}\neq Y_{\leq n-1})\cap{\texttt{AccurateN}_{\leq n-1}})$
+Observe that $\overline{\texttt{AccurateN}_{\leq n-1}}=\bigcup_{\ell=1}^{n-1}\overline{\texttt{AccurateN}_{\ell}}$.
+Therefore, we can further applying the above recurrence, in particular, for every $\ell$, we will use the following upper bound.
+$\overline{\texttt{AccurateN}_{\ell}}\subseteq$ $(\overline{\texttt{AccurateN}_{\ell}}\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}))\cup(X_{\leq\ell-1}\neq Y_{\leq\ell-1})$
+Therefore, $\quad\overline{\texttt{AccurateN}_{n}}\subseteqTherefore$ $\bigcup_{\ell=1}^{n}(\overline{\texttt{AccurateN}_{\ell}}\cap{\texttt{AccurateN}_{\leq\ell-1}}\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}))$
+$\qquad\qquad\cup\bigcup_{\ell=1}^{n-1}((X_{\leq\ell-1}\neq Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell-1}})$
+Let us focus on two successive terms of the second expression and using the equation $A=A\cup(\overline{A}\cap B)$
+$\underbrace{\left((X_{\leq\ell-1}\neq Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell-1}}\right)}_{A}\cup\underbrace{\left((X_{\leq\ell}\neq Y_{\leq\ell})\cap{\texttt{AccurateN}_{\leq\ell}}\right)}_{B}$
+$=$ $((X_{\leq\ell-1}\neq Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell-1}})\cup((X_{\leq\ell-1}=Y_{\leq\ell-1})\cap(X_{\leq\ell}\neq Y_{\leq\ell})\cap{\texttt{AccurateN}_{\leq\ell}})$
+$\quad\quad\cup(\overline{\texttt{AccurateN}_{\leq\ell-1}}\cap(X_{\leq\ell}\neq Y_{\leq\ell})\cap{\texttt{AccurateN}_{\leq\ell}})$
+Noting that $((X_{\leq\ell-1}=Y_{\leq\ell-1})\cap(X_{\leq\ell}\neq Y_{\leq\ell})=((X_{\leq\ell-1}=Y_{\leq\ell-1})\cap(X_{\ell}\neq Y_{\ell})$ and the fact that $(\overline{\texttt{AccurateN}_{\leq\ell-1}}\cap{\texttt{AccurateN}_{\leq\ell}})=\varnothing$ we have,
+$((X_{\leq\ell-1}\neq Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell-1}}\cup((X_{\leq\ell}\neq Y_{\leq\ell})\cap{\texttt{AccurateN}_{\leq\ell}})$
+$=$ $((X_{\leq\ell-1}\neq Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell-1}})\cup\left((X_{\ell}\neq Y_{\ell})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell}}\right)$
+Applying the above simplification for all terms,
+$\bigcup_{\ell=1}^{n-1}((X_{\leq\ell-1}\neq Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell-1}})=\bigcup_{\ell=1}^{n-1}\left((X_{\ell}\neq Y_{\ell})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell}}\right)$
+[…]
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\cup\bigcup_{\ell=1}^{n-1}\left((X_{\ell}\neq Y_{\ell})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell}}\right)$